### v1

# On the Effects of Randomness on Stability of Learning with Limited Labelled Data: A Systematic Literature Review

Abstract.
Learning with limited labelled data, such as few-shot learning, meta-learning or transfer learning, aims to effectively train a model using only small amount of labelled samples. However, these approaches were observed to be excessively sensitive to the effects of uncontrolled randomness caused by non-determinism in the training process. The randomness negatively affects the stability of the models, leading to large variance in results across training runs. When such instability is disregarded, it can unintentionally, but unfortunately also intentionally, create an imaginary perception of research progress. Recently, this area started to attract a research attention and the number of relevant studies is continuously growing. In this survey, we provide a comprehensive overview of 134 papers addressing the effects of randomness on the stability of learning with limited labelled data. We distinguish between four main tasks addressed in the papers (investigate/evaluate; determine; mitigate; benchmark/compare/report randomness effects), providing findings for each one. Furthermore, we identify and discuss seven challenges and open problems together with possible directions to facilitate further research. The ultimate goal of this survey is to emphasise the importance of this growing research area, which so far has not received appropriate level of attention.
randomness, stability, meta-learning, transfer learning, language model fine-tuning, in-context learning, literature survey
ccs: General and reference Surveys and overviewsccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Machine learningccs: Computing methodologies Learning paradigms
1. Introduction
The approaches for learning with limited labelled data are designed to achieve high performance of the machine learning models even with few labels available (Brown et al., 2020; Wang et al., 2020, 2019c; Song et al., 2023). Under the term learning with limited labelled data we understand any approach that is designed to work with the lack of labels, without any constraint on how the labelled samples are distributed, i.e., whether all the samples are from a single task or are distributed across different tasks. Although similar to the notion of few-shot learning, it represents a broader scope encompassing a larger number of possible approaches (some of which are incorrectly categorised as few-shot learning in current practice).
To deal with the limited labels, the approaches utilise additional information from different sources (Brown et al., 2020; Wang et al., 2020, 2019c). A popular option is to transfer knowledge from similar tasks and datasets using transfer learning. For example, in the NLP domain it is common to use language model fine-tuning, where a language model pre-trained on a large set of unlabelled data (from unrelated domain) is fine-tuned for the specific downstream task or problem statement using only few labelled samples available for it (Xia et al., 2020; Rogers et al., 2020; Mosbach et al., 2021; Dodge et al., 2020). Another emerging technique is the in-context learning, where a large pre-trained language model is “prompted“ to predict a label for a single test sample by presenting it with a concatenation of few labelled samples and the test sample, without requiring any parameter update (Brown et al., 2020; Liu et al., 2023). Finally, meta-learning is often used, where the model is explicitly trained to quickly adapt to a new task with only a handful of examples by learning how best to learn across large number of related tasks with few labelled samples each (Agarwal et al., 2021; Hospedales et al., 2021; Kadam and Vaidya, 2018).
However, a significant problem observed for the approaches for learning with limited labelled data is their sensitivity to the effects of uncontrolled randomness, which negatively affects their stability. Under stability, we understand a property of a learning algorithm or a model that indicates what influence the small scale and random changes (or perturbations) in the input data and parameters have on its outputs (Liu et al., 2017). Such random changes are introduced by different randomness factors that represent the non-deterministic decisions in the training process, such as random initialisation of model parameters, or data shuffling (Pham et al., 2021; Gundersen et al., 2022). These randomness factors represents the main point around which the effects of randomness on the stability are addressed.
The uncontrolled effects of randomness can have massive impact on the stability of the utilised machine learning approaches. Running the training multiple times on the same dataset, with the same setup and hyperparameters may lead to large deviation in the final performance (Mosbach et al., 2021; Dodge et al., 2020; Lu et al., 2022; Agarwal et al., 2021). Repeating BERT fine-tuning multiple times and cherry-picking the best result, it is possible to outperform more recent BERT variants (Dodge et al., 2020). Changing only order of data in in-context learning can lead the model from state-of-the-art predictions to random guesses (Lu et al., 2022). Or choosing different set of adaptation data in meta-learning can lead to difference between minimum and maximum performance being up to 90% (Agarwal et al., 2021). Even though the effects of randomness are present in all cases of machine learning, their impact on the stability is especially significant in low data regimes.
If the uncontrolled randomness is not appropriately addressed in such low data regimes, it can have a significant non-negligible negative consequences. In research works and benchmarks, it may prohibit an objective performance comparison of newly designed methods to the state-of-the-art baselines (especially when the difference between their performance is similar to the deviation caused by the randomness). In such case, a method can be incorrectly denoted as state-of-the-art only based on a more favourable random chance (Reimers and Gurevych, 2017) (as illustrated in Figure 0(b)). The uncontrolled randomness can unintentionally, but unfortunately also intentionally (by cherry-picking), create an imaginary perception of research progress.
The effects of randomness are addressed in various depths - from a pure recognition of randomness effects, through a deeper investigation up to a full mitigation of such effects (illustrated in the Figure 1). Recent few-shot learning surveys recognise the effects of randomness as an open problem (Wang et al., 2020; Tian et al., 2022; Dong et al., 2022), or even dedicate a part of the survey to a basic overview of these effects (Song et al., 2023). Other papers simply recognise the impact of randomness on the overall results and explicitly repeat the training and evaluation multiple times to obtain a more representative, albeit aggregated value to address the randomness. However, no further analysis regarding the effects of the randomness is provided. Finally, only a small fraction of papers provide more detailed analysis of the effects of randomness to improve the stability and reduce the significant deviation in results. These works conduct a series of experiments that are designed to investigate the effects, estimate the distribution of results and determine the significance and the impact of the different randomness factors on the stability. The effects of the randomness are analysed in more detail, in order to determine the real origin of randomness in the training process (e.g., under-specification, which causes slightly different parameter initialisation to converge to different minima). Based on the findings from the analysis of the experiments, various methods are proposed to mitigate the effects of randomness, improving the stability and reducing the deviation to achieve more stable results.
(a) Various depths the effects of randomness can be addressed in the papers
(b) Comparison based on a single value and based on distribution of the results, inspired by (Reimers and Gurevych, 2017)
Figure 1. The effects of randomness are addressed in various depths in the papers. First, the effects are simply recognised, resulting in reporting of a single aggregated value from multiple repeated training and evaluation runs (illustrated as “Recognise”). This leads to biased comparison between approaches, as one approach may show better results only due to random chance and unintentional cherry-picking (illustrated in Figure 0(b)). Analysing the effects of randomness, estimating the distribution of results from multiple runs and comparing the distributions can prevent this problem (illustrated as “Describe”). As the distribution is not affected in any way, only estimated, the deviation is not reduced. The effects of randomness are properly addressed when they are mitigated, reducing the deviation in the distribution of results (illustrated as “Mitigate”). However, an effective mitigation requires understanding of the effects, their importance, and the origin of randomness provided by the analysis of the effects of randomness (provided by the “Describe” task).
In this paper, we provide a comprehensive survey of 134 papers that address the effects of randomness, either as simply recognising the effects of randomness on the stability (N=84), or investigating and describing the characteristics of these effects in more detail and/or exploring the appropriate means of their mitigation (N=50). We provide a more in-depth analysis of the 50 papers that address the effects of randomness in more detail. We provide aggregated findings from this analysis and use them to identify seven open problems related to the study of the effects of randomness, while providing future directions on how to address them.
This is the first survey of its kind that provides a comprehensive and systematic literature review specifically focused on the impact of the effects of randomness on stability across various approaches for dealing with limited labelled data. As opposed to previous papers that provide an overview of possible randomness factors and investigate these effects (Pham et al., 2021; Gundersen et al., 2022), we focus on all tasks for addressing the effects of randomness, not only their investigation, but also determining their origin, mitigation and benchmarking in presence of these effects of randomness.
The purpose of this survey is to emphasise the importance of this research area, which so far has not received appropriate level of attention from researchers and practitioners. At first, it should serve existing or new-coming researchers who are explicitly tackling with the topic of the randomness and its effects on the stability of learning with limited labelled data to support their research (by providing the overview of the state-of-the-art, identification of open problems and future direction). Secondly, its purpose is also to inform researchers and practitioners utilising the learning with limited labelled data about the consequences of unaddressed randomness and how to effectively prevent them. To better achieve these purposes, we summarised all the identified papers along with their categorisation as in the supplementary material of this survey (Section A).
The rest of the paper is structured as follows. In Section 2, we describe the scope of the survey, difference to existing surveys and the methodology used for identifying relevant papers. Taxonomy applied to analyse and categorise the identified papers is described in Section 3. Afterwards, we describe the different tasks for addressing the effects of randomness, categorised based on our taxonomy, in Sections 4, 5, 6 and 7, with a summary of the findings at the end of each such section. The identified challenges and open problems are described in Section 8. Finally, the survey is concluded in Section 9 with summary of our contributions and findings.
2. Background
2.1. Scope of the Survey
In this survey, we focus on a specific part of stability. We focus on papers that explore the effects of small perturbations to multiple inputs of the training process on the final performance after the training process is done (as opposed to the effects of the perturbation of a single sample on the loss or gradient of training as defined by uniform stability (CHEN et al., 2020)). A similar, but at the same time quite distinct property is the robustness of the model. It deals with a more large scale or systematic changes in the input, such as distribution shift or presence of adversarial examples. As the adversarial robustness is already extensively studied in the context of learning with limited labelled (e.g., in meta-learning (Goldblum et al., 2020; Yin et al., 2018; Xu et al., 2021a; Wang et al., 2021b)) and the effects are often systematic, we consider it to be out of scope for this survey. Finally, we view the stability to be a subset of the reproducibility of the model. Besides randomness in the training process, the reproducibility is impacted by other factors such as selective reporting, biases, data and source code availability, and many others as specified in (Gundersen et al., 2022). These additional factors are out of scope for this survey.
In case of machine learning approaches, we mainly focus on the approaches that utilise related tasks and datasets as additional information. Not requiring large set of labelled data, these approaches are usable across broader set of domains, making them also more popular. Due to this popularity, they also cover almost the whole research area of addressing the effects of randomness on the stability, as there are almost no other works focusing on approaches that utilise information from the unlabelled sets of data. Our focus is on three prevalent groups of approaches: 1) meta-learning; 2) language model fine-tuning; and 3) in-context learning. Even though we do not explicitly focus on a specific modality, the popularity of the approaches impact the scope of this survey. As the majority of the papers that focus on stability of transfer learning work with language models, we define the scope in this survey using the language model fine-tuning and in-context learning approaches for better readability. We provide only the general problem statement for each group of approaches (in Section 3.3), while providing references to more comprehensive overview.
2.2. Existing Surveys
To the best of our knowledge, there are no existing surveys focused on addressing the effects of randomness on stability. However, there are surveys on related topics that partially touch on the topics of stability. First of all, some papers provides an overview of possible sources of randomness, although with focus on replicability (Gundersen et al., 2022). In some surveys on the topic of learning with limited labelled data and/or few-shot learning, the stability and the effects of randomness are mentioned as a problem that should be addressed, with varying level of detail provided (Li et al., 2021; Tian et al., 2022; Dong et al., 2022; Song et al., 2023).
At the same time, the study of effects of randomness on the stability is becoming a research topic addressed in an increasing number of research works focused on dealing with the lack of labels, mainly focused on the use of pre-trained models (Dodge et al., 2020; Mosbach et al., 2021; Lu et al., 2022). Even first benchmarks consider the stability of the models to be paramount and include it as a part of the final metric (Efrat et al., 2022).
To fill this gap, in this survey we aim to provide the first comprehensive overview of the effects of randomness on the stability of learning with limited labelled data and describe it from perspectives that are not easily visible from the standard analysis of related works.
Figure 2. Process for identifying and categorising papers for this survey. Due to strong clustering effects, additional papers are identified using a reference analysis on the most relevant identified papers, i.e., including papers cited in them, as well as the ones that cite them.
2.3. Survey Methodology
To identify the set of relevant papers for our comprehensive survey, we take inspiration from the PRISMA methodology (Moher et al., 2009; Page et al., 2021), performing following steps (as illustrated in Figure 2):
  1. (1)
    Identification of relevant papers. A keyword search query is defined based on the related terms for our defined scope identified in different papers. Multiple digital libraries are searched using the defined query to identify set of potentially relevant papers.
  2. (2)
    Relevance filtering. The potentially relevant papers are categorised into irrelevant and relevant based on the context the defined keywords are used in.
  3. (3)
    Content filtering. The non-full papers or papers released in unrelated venues, such as extended abstracts, proposals for talks or papers appearing in unrelated workshops not dealing with stability, are filtered out.
  4. (4)
    Paper analysis. The papers are analysed based on their depth of addressing the randomness effects. The papers that only recognise the problem of stability without further investigation are referred to as recognise papers. The remaining papers are referred to as core papers.
  5. (5)
    Reference analysis. In order to discover additional potentially relevant papers, the papers cited in the core papers; as well as papers citing the core papers, are examined and their relevance is determined based on their title and abstract. The papers deemed potentially relevant are an additional input into the second step of this process.
  6. (6)
    Analysis and categorisation. The identified core papers are analysed and categorised based on the taxonomy (as defined in Section 3) to provide the comprehensive overview and to identify challenges and open problems.
As a result, 134 papers are identified — out of them, 84 papers are classified as recognise papers, and 50 as core papers. The number of identified papers grouped by the publication year is presented in Figure 3. The categorisation of the core papers is showcased in Table 1. Moreover, the full list of all identified papers (including recognise papers), with their categorisation and additional metadata is available in the supplementary material (Section A). For a more detailed implementation of the survey methodology, see the supplementary material (Section C).
(a) Core papers only
(b) All papers
Figure 3. Number of papers dealing with randomness grouped by the year. Figure 2(a) shows only the core papers that focus on addressing the effects of randomness in more detail, while Figure 2(b) also includes the papers that only recognise the problem. The problem started to attract attention only in recent years.
Table 1. A showcase of the core papers and their categorisation based on the proposed txonomy. The full categorisation, along with the recognise papers is available in the supplementary material (Section A).
References Area Factor Performed Tasks Modality Interactions Out-Of-Distribution
Machine Learning
Meta-Learning
LM Fine-Tuning
In-Context Learning
Label Selection
Data Split
Task Choice
Data Choice
Order of Data
Parameter Initialisation
Random Seed
Implementation
Number of Samples
Hyperparameters
Augmentation
Noise
Approach Specific
Investigate + Evaluate
Determine
Mitigate
Benchmark + Compare + Report
Text
Image
(Bouthillier et al., 2021) $\times$
(Agarwal et al., 2021) $\times$ $\times$
(Boquet et al., 2019) $\times$ $\times$
(Lu et al., 2022) - $\times$ $\times$
(Zhao et al., 2021b) $\times$
(Osahor and Nasrabadi, 2022) $\times$ $\times$
(Zhang et al., 2021a) $\times$ $\times$
(Min et al., 2022) - $\times$ $\times$
(Chen et al., 2022) $\times$ $\times$
(Winata et al., 2021) - $\times$ $\times$
(Chang et al., 2021) $\times$ $\times$
(Schick and Schütze, 2021) - $\times$ $\times$
(Jundi and Lapesa, 2022) $\times$
(Wu et al., 2022) $\times$ $\times$
(Yang and Kwok, 2022) $\times$ $\times$
(Zhao et al., 2021a) - $\times$
(Setlur et al., 2021b) $\times$
(Dror et al., 2019) $\times$ $\times$
(Dodge et al., 2020) $\times$
(Setlur et al., 2021a) $\times$ $\times$
(Mosbach et al., 2021) $\times$ $\times$
(Kumar and Talukdar, 2021) - $\times$ $\times$
(Reimers and Gurevych, 2017) $\times$ $\times$
(Zhang et al., 2021b) $\times$
(Liu et al., 2022) - $\times$ $\times$
(Sellam et al., 2022) $\times$
(McCoy et al., 2020) $\times$
(Zhou et al., 2020) $\times$
(Bouthillier et al., 2019) $\times$ $\times$
(Dehghani et al., 2021) $\times$ $\times$
(Zhong et al., 2021) $\times$
(Dauphin and Schoenholz, 2019) $\times$ $\times$
(D’Amour et al., 2022) $\times$ $\times$
(Khurana et al., 2021) $\times$ $\times$
(Hua et al., 2021) $\times$ $\times$
(Hidey et al., 2022) $\times$ $\times$
(Pham et al., 2021) $\times$ $\times$
(Cioba et al., 2022) $\times$
(Phang et al., 2019) $\times$
(Wang et al., 2021a) $\times$ $\times$
(Zhang et al., 2022b) $\times$
(Lee et al., 2020) $\times$ $\times$
(Xu et al., 2021c) $\times$ $\times$
(Meng et al., 2022) $\times$ $\times$
(Zhang et al., 2022a) - $\times$
(Zheng et al., 2022) $\times$ $\times$
(Ye et al., 2021) $\times$
(Bragg et al., 2021) $\times$ $\times$
(Mukherjee et al., 2021) $\times$ $\times$
(Efrat et al., 2022) - $\times$ $\times$
3. Taxonomy for Literature Analysis and Categorisation
The identified core papers are categorised based on multiple dimensions that describe their characteristics of interest. We use two groups of characteristics: 1) primary categorisation characteristics, which provide the primary means for dividing and categorising the individual papers; and 2) secondary categorisation characteristics, which provide a more nuanced clustering of papers in the specific primary categorisation dimensions.
We define three separate primary properties that are used to categorise the papers: 1) what tasks are performed to address the randomness; 2) randomness factors addressed by the papers; and 3) what machine learning approach the papers use. The performed tasks property is the main property based on which we organise the rest of the survey, while the randomness factors and machine learning approach properties are mostly used to define the scope in this survey (see Section 2.1). We provide the paper distribution across individual randomness factors and specific machine learning approaches for all core papers in Table 2 (the paper distributions for specific tasks are included in the supplementary material, Section A).
Table 2. Relation between the randomness factors and machine learning approach across all analysed papers. We can observe different popularity of the randomness factors across different machine learning approaches. The majority focus being denoted in bold. For example, fine-tuning mostly focuses on the random seeds.
All Meta-Learning Fine-Tuning In-Context Learning
(N=8) (N=10) (N=24) (N=13)
Label Selection 1 2 2 1
Data Split 2 2 1
Task Choice 1 5 N/A
Data Choice 1 6 2 6
Random Seed 4 2 17 5
Model Initialisation 5 8 N/A
Order of Data 4 1 7 9
Implementation 1 1
Number of Samples 5 6
Hyperparameters 3 1 4 1
Augmentation 1
Noise 1 1 1
Approach Specific N/A 2 3
3.1. Tasks Performed to Address the Randomness
This property specifies the different tasks for addressing the effects of randomness that are performed in the papers (the relations between these tasks, along with their input and output are illustrated in Figure 4). For this property, we define following 7 values:
  1. (1)
    Investigate - the effects of randomness from different randomness factors are investigated using specifically designed experiments. For example, the training and evaluation is done multiple times to determine the distribution of performance across multiple runs.
  2. (2)
    Determine - an effort to discover the origin of the randomness (different from the randomness factors) is made by analysing the behaviour of the specific machine learning approach or model in presence of randomness. For example, to determine why changing the order of samples causes significant change in performance, the experiments that observe specific problems (e.g., label bias) are designed and analysed.
  3. (3)
    Mitigate - mitigation strategies for reducing the effects of randomness are proposed, used and their effectiveness is determined. For example, it is proposed to use more deterministic data sampling based on heuristics instead of random sampling.
  4. (4)
    Benchmark - a benchmark that is specifically designed to take into consideration the effects of randomness is proposed. For example, a benchmark that performs multiple train and test splits and evaluates on all of them.
  5. (5)
    Evaluate - a more sophisticated, usually statistical, framework for evaluating the experiments investigating the effects of randomness, or determining the effectiveness of mitigation strategies, is proposed. For example, instead of calculating a simple average value, it is proposed to estimate the distribution of results and evaluate the difference using statistical tests.
  6. (6)
    Report - it is proposed to report more than a single value when running and evaluating experiments and comparing approaches to better deal with the effects of randomness. Not necessarily dependent on experiments for addressing the effects of randomness, but also recommendations how to improve the reporting of results overall when the randomness has a strong effect on performance. For example, instead of reporting single average value, the minimum and maximum value over multiple runs should be reported as well.
  7. (7)
    Compare - a better comparison strategies between different approaches in presence of significant effects of randomness are proposed. For example, instead of comparing single average values, it is proposed to compare distributions of results using statistical tests.
We aggregate the closely related tasks together and organise the survey based on these aggregated groups. Namely, we aggregate together the investigate and evaluate tasks, as the evaluation strategies are designed specifically for the investigation experiments. In addition, we aggregate the benchmark, compare and report values, as they are closely related.
As the different papers can focus on different, often disjointed, tasks for addressing the effects of randomness, this property does not necessarily provide a distinct division of the papers. For example, the most common combination is to perform investigation, determination and mitigation of a specific randomness factor in a single paper. In such cases, the paper is mentioned in this survey multiple times, each time focusing only on the parts relevant for the specific task.
Figure 4. Different tasks for addressing the effects of randomness, along with their inputs, outputs and the relations between them. The dashed lines represent relations between the tasks that currently do not exist, but need to be considered in the optimal state of addressing the effects of randomness, e.g., using more sophisticated evaluation when determining the effectiveness of mitigation strategies. The related tasks are grouped together, e.g., evaluate task near the investigate task.
3.2. Randomness Factors
The randomness factors property specifies what randomness factors are actually addressed in the paper in some way (i.e., their effects are investigated or mitigated), but not factors that are just mentioned (e.g., mention that other papers found the factor to be unimportant). Papers can focus on multiple factors at the same time.
Multiple randomness factors have already been identified and described in the existing works (e.g., (Pham et al., 2021; Gundersen et al., 2022)). Since the used terminology differs across the works, for the purpose of this survey, we established the grounds (in terms of common terminology and understanding) by collecting these randomness factors and unifying their denomination and definition.
The identified randomness factors can be categorised into three distinct groups: 1) factors related to the input data; 2) factors related to parameters of the model and the overall training process; and 3) factors related to the underlying implementation of deep learning frameworks and hardware.
As the randomness factors related to the input data, we identify following:
  1. (1)
    Label selection - specifies what data is considered to be annotated from larger set of unlabelled data. Represents the real world scenario with limited labelled budget, where only specific samples are annotated. Has potentially significant impact on the underlying data distribution.
  2. (2)
    Data split - randomness introduced by different splits of data into training, validation and testing set. Mainly impacts the information provided to model during training.
  3. (3)
    Task choice - randomness introduced by choosing different tasks for training. This factor is mainly relevant for approaches that perform few-shot learning, such as meta-learning, or when using a benchmark dataset composed of different tasks.
  4. (4)
    Data choice - randomness introduced by the sampling of data the model uses for adaptation (meta-learning), or as examples in prompts (in-context learning).
  5. (5)
    Order of data - randomness introduced by the random shuffling of data in different training iterations. Mainly relevant for the in-context learning (order of labelled samples in the prompt) and fine-tuning. For meta-learning, this factor represents the order in which the different tasks are presented during training.
  6. (6)
    Augmentation - randomness introduced by different data augmentation strategies.
  7. (7)
    Noise - randomness introduced by adding noise to the different samples. In essence, it can be viewed as a subset of the augmentation randomness factor.
  8. (8)
    Number of samples - specifies the randomness introduced by choosing different number of labelled samples in training, such as different number of examples in in-context learning. It can be viewed as an edge case with regards to the definition of stability, as the expected change is often systematic (but is not always the case).
Similarly, we identify the following randomness factors corresponding to the randomness related to model parameters and the overall training process:
  1. (1)
    Parameter initialisation - randomness introduced by randomly initialising the parameters of the model, such as weights in neural networks. Mainly relevant for meta-learning and language model fine-tuning, as no initialisation is performed in in-context learning.
  2. (2)
    Random seed - randomness introduced by non-deterministic actions in the training process, such as the use of non-deterministic layers. Often also used as an aggregated factor for the parameter initialisation and the order of data randomness factors.
  3. (3)
    Hyperparameters - randomness caused by introducing small random changes into hyperparameters. Can also be used when performing hyperparameter optimisation in the low data regimes, due to large effects of randomness on them. Also aggregates the randomness introduced by changing parameters related to optimisation such as parameters used in optimiser or the number of training iterations.
  4. (4)
    Approach specific parameters - aggregates the randomness introduced by factors specific for only some of the approaches. Includes factors like different prompt design (in-context learning), or different task sampling strategies (meta-learning).
Randomness introduced by implementation of deep learning frameworks and hardware is a special low level group, which includes factors such as scheduling and floating point precision, auto-selection of primitive operations, how parallel processes are used, changes introduced by different implementation of frameworks such as TensorFlow or PyTorch and their different software versions, but also many other factors influencing reproducibility more than the stability of results (Pham et al., 2021; Gundersen et al., 2022). Since these randomness factors cannot be easily addressed, for the purpose of this survey, we do not specifically distinguish between them and consider them jointly as one group.
Each randomness factor is associated with a set of its randomness factor configurations (sometimes referenced in short as configurations). A single configuration specifies one state from the set of all possible states the randomness factor can attain. For example, a single randomness factor configuration for the order of data specifies one possible permutation of the samples, a single set of weights used for initialising the whole model, or a single random value to seed the random number generator. Using this definition, the set of randomness factor configuration can be infinite for some factors (e.g., initialisation weights) or finite, but extremely large (e.g., permutations of order).
The different randomness factors may depend on each other, leading to interactions between them. An example of such interactions is the relation between data choice and order of data randomness factors, where the choice of what data we use in training explicitly affects the order of samples. Therefore, such interactions can lead to negative, often unforeseen consequences, such as prohibiting reliable freezing of randomness factor configurations (e.g., it is not possible to have fixed order of samples when using different choices of data).
3.3. Machine Learning Approach
Similarly, the machine learning approach property specifies which group of approaches is addressed in the paper. Besides the machine learning approaches that define the scope of this survey (Section 2.1), we also consider a generic machine learning. This is a special value that contains all papers that still deal with effects of randomness, but do not utilise any of the meta-learning, fine-tuning or in-context learning approaches. The main idea, high-level overview and reference to popular approaches and comprehensive surveys of the machine learning approaches we work in this survey are included in the supplementary material (Section B).
3.4. Secondary Categorisation Properties
Following additional properties are used for more nuanced categorisation and clustering of papers:
  1. (1)
    Modality that divides the papers into those focusing on text data, images, tabular data, or any combination of these modalities.
  2. (2)
    Groups of approaches and models further divide the primary property of the machine learning approach into a more detailed categorisation. For example, the meta-learning approaches are further divided into optimisation or metric based meta-learning, and then even further into the specific approaches like MAML (Finn et al., 2017), or Prototypical Networks (Snell et al., 2017).
  3. (3)
    Interactions, a binary property that specifies whether the confounding effects between different randomness factors are addressed in any way in the paper.
  4. (4)
    Standard deviation, a binary property that specifies whether the performance is reported using standard deviation (or similar metric such as confidence intervals) or just as a simple aggregate metric (mean, median, etc.).
  5. (5)
    Out-of-distribution, a binary property that specifies whether the data used for testing purposes in the papers comes from the same distribution as the training data, such as addressing the effects of randomness when dealing with multilingual data or data drifts.
  6. (6)
    Datasets used to better explore what datasets are popular when addressing the effects of randomness. We do not list all the possible values as the list is extensive, but includes popular datasets such as GLUE (and its sub-datasets) (Wang et al., 2019b, a), MiniImageNet (Vinyals et al., 2016) or CIFAR-FS (Bertinetto et al., 2019).
4. Investigating the problem of stability
In this task, the effects of randomness on the stability are investigated, their impact is estimated and the significance of this impact is determined. Each paper investigating the effects of randomness can be characterised by three main aspects: 1) scope of the investigation; 2) investigation methodology; and 3) results from the investigation. The scope of the investigation includes decisions about what factors to investigate or across what models, datasets and modality the investigation is performed. It also includes more nuanced specifics like what problem statement is considered (simple classification/regression or sequence generation), or from what setting the data comes from (in-distribution vs. out-of-distribution). Although these decisions have minimal impact on the methodology, they strongly govern the obtain results and their generality.
The investigation methodology determines how the investigation is done. It provides answers to different problems and specifics of the investigation, such as how to setup hyperparameters, how to perform the investigation (multiple fully randomised runs or keeping some parts static), how many runs to use, how to address interactions between randomness factors or how to evaluate the different runs. Again, each decision has significant impact on the results of the investigation, mainly how trustworthy and objective they are.
Finally, the results of the investigation specify findings and the behaviour that was observed. Interpreting these results is important for informing the further tasks for addressing the effects of randomness (i.e., determining real origin of randomness and mitigating the effects). They can also be used to provide recommendations on the best use of models in practice in low data regimes.
4.1. Scope of the investigation
Majority of the papers focus on investigating how the randomness affects the language models (14 out of 24) when they are used to deal with limited labelled data. The focus is distributed evenly across both the fine-tuning of language models (Sellam et al., 2022; Zhao et al., 2021b; Jundi and Lapesa, 2022; Dodge et al., 2020; McCoy et al., 2020; Zhou et al., 2020; Zhong et al., 2021) and the use of in-context learning (Lu et al., 2022; Winata et al., 2021; Schick and Schütze, 2021; Zhao et al., 2021a; Liu et al., 2022; Zhang et al., 2022a; Efrat et al., 2022). The focus on investigating overall machine learning approaches (mainly their use with few data and evaluation) (Bouthillier et al., 2021; Reimers and Gurevych, 2017; Bouthillier et al., 2019; Dehghani et al., 2021; Pham et al., 2021) and the use of meta-learning (Boquet et al., 2019; Agarwal et al., 2021; Setlur et al., 2021a; Cioba et al., 2022; Ye et al., 2021) is lower, but not by a high margin.
The significant focus on language models also translates to what modality is addressed the most. Investigating text data in NLP problems gets the most focus (Bouthillier et al., 2021; Sellam et al., 2022; Lu et al., 2022; Zhao et al., 2021b; Winata et al., 2021; Schick and Schütze, 2021; Jundi and Lapesa, 2022; Zhao et al., 2021a; Dodge et al., 2020; Reimers and Gurevych, 2017; Liu et al., 2022; McCoy et al., 2020; Zhou et al., 2020; Zhong et al., 2021; Zhang et al., 2022a; Ye et al., 2021; Efrat et al., 2022), with images being mostly addressed only in meta-learning and machine learning approaches (but not all of them) (Boquet et al., 2019; Agarwal et al., 2021; Setlur et al., 2021a; Bouthillier et al., 2019; Pham et al., 2021; Cioba et al., 2022), with only a single paper considering investigation across both image and text data (Dehghani et al., 2021).
Some randomness factors receive more attention in investigation than other ones. The more popular factors are usually the ones that are relevant for all approaches. The randomness factor that receives the most attention is the random seed, which aggregates both the initialisation and the order of data in a single value. The papers can be separated into those that consider only random seed (Boquet et al., 2019; Zhao et al., 2021b; Schick and Schütze, 2021; Reimers and Gurevych, 2017; Zhou et al., 2020; Dehghani et al., 2021; Ye et al., 2021) and the ones that explicitly specify all the factors affected by changing the configuration of random seed configuration (Sellam et al., 2022; Dodge et al., 2020; Zhong et al., 2021; Bouthillier et al., 2021; McCoy et al., 2020; Bouthillier et al., 2019; Pham et al., 2021). Some papers consider only the initialisation (Jundi and Lapesa, 2022; Liu et al., 2022) or order of data (Lu et al., 2022; Winata et al., 2021; Zhao et al., 2021a; Zhang et al., 2022a; Efrat et al., 2022) as single factor without the combination with the other one. Another quite popular general randomness factor is the data and task choice (Sellam et al., 2022; Agarwal et al., 2021; Zhao et al., 2021b; Schick and Schütze, 2021; Zhao et al., 2021a; Setlur et al., 2021a; Liu et al., 2022; Dehghani et al., 2021; Zhang et al., 2022a; Ye et al., 2021; Efrat et al., 2022). Other factors that influence all approaches, such as effects of augmentation (Bouthillier et al., 2021), the choice of hyperparameters (Bouthillier et al., 2021; Boquet et al., 2019; Reimers and Gurevych, 2017), choice of what data is labelled (Bouthillier et al., 2021; Jundi and Lapesa, 2022; Setlur et al., 2021a; Liu et al., 2022; Cioba et al., 2022), split of data (Bouthillier et al., 2021; Dehghani et al., 2021), noise (Bouthillier et al., 2021) and implementation (Boquet et al., 2019; Pham et al., 2021) are getting less attention. There is also little focus on the more niche randomness factors such as prompt design (Efrat et al., 2022; Zhao et al., 2021a), early stopping (Dodge et al., 2020) or the number of samples used (Lu et al., 2022; Zhao et al., 2021b, a; Jundi and Lapesa, 2022; Liu et al., 2022). However, the popularity and focus on the individual randomness factors is dependent on the approaches for learning with limited labelled data and their popularity, as some randomness factors are solely investigated only in specific approaches and not across all of them.
The choice of what datasets and what models are used for investigation influences how general the final results are. Many papers consider only a single dataset or a single model architecture in their investigation (Pham et al., 2021; Setlur et al., 2021a; McCoy et al., 2020; Winata et al., 2021; Bouthillier et al., 2021; Dodge et al., 2020; Sellam et al., 2022), limiting the generalisability of the results. It is more common to investigate the effects of randomness across different models, either different architectures, or the same architecture with different numbers of parameters, and across different datasets (Bouthillier et al., 2019; Agarwal et al., 2021; Cioba et al., 2022; Ye et al., 2021; Zhou et al., 2020; Zhao et al., 2021b; Jundi and Lapesa, 2022; Schick and Schütze, 2021; Liu et al., 2022; Zhao et al., 2021a; Zhang et al., 2022a; Boquet et al., 2019). In some specific cases, the investigation is specifically designed as an investigation of the randomness across models of different sizes, or as a comparison between different model architectures (Reimers and Gurevych, 2017; Dehghani et al., 2021; Zhong et al., 2021; Lu et al., 2022; Efrat et al., 2022), mainly to investigate how the effects of randomness affect the choice of what model to use.
In the majority of the papers, the investigation is performed in an in-distribution setting.Only few papers consider the investigation in out-of-distribution setting to be important (McCoy et al., 2020; Zhou et al., 2020; Cioba et al., 2022; Ye et al., 2021), often masked in cross-lingual problem statement.
4.2. Investigation Methodology
Overall, the differences in how the investigation of the effects of different randomness factors is done is small across different papers that focus on different approaches and datasets, making the investigation quite consistent. In all papers, the investigation is performed by changing the randomness factor configuration (such as changing the value of random seed) and observing the changes in behaviour of the model and its results. The most common way is to introduce the change randomly into the configuration (Reimers and Gurevych, 2017; Bouthillier et al., 2019; Cioba et al., 2022; Ye et al., 2021; McCoy et al., 2020; Zhou et al., 2020; Zhao et al., 2021b; Jundi and Lapesa, 2022; Zhong et al., 2021; Schick and Schütze, 2021; Liu et al., 2022; Zhang et al., 2022a; Bouthillier et al., 2021; Sellam et al., 2022). To exert additional control over the investigation and its results, mainly when considering multiple randomness factors at the same time and addressing the interactions between them, some specific randomness factor configurations are fixed (usually the ones not investigated), while randomly varying the rest (usually the ones that we want to investigate) (Pham et al., 2021; Lu et al., 2022; Zhao et al., 2021a; Efrat et al., 2022; Boquet et al., 2019; Dodge et al., 2020). A combination of previous two approaches can also be used to compare between them (Winata et al., 2021). In specific cases when the number of possible configurations is small, the whole space of possible configurations is explored, such as trying all the possible permutations when investigating the effects of the order of data with few samples (Dehghani et al., 2021; Agarwal et al., 2021). Finally, an importance of a specific randomness factor configuration is determined by keeping it fixed throughout the whole investigation (Setlur et al., 2021a).
An important aspect of the investigation is how the hyperparameters are handled. Majority of papers use default hyperparameters or find an optimal set for each model and dataset and use it throughout the whole investigation (Reimers and Gurevych, 2017; Dehghani et al., 2021; Pham et al., 2021; Agarwal et al., 2021; Setlur et al., 2021a; Cioba et al., 2022; Ye et al., 2021; Zhou et al., 2020; Zhao et al., 2021b; Jundi and Lapesa, 2022; Zhong et al., 2021; Boquet et al., 2019), or reuse a fixed set of hyperparameters determined to be optimal in previous works (McCoy et al., 2020; Dodge et al., 2020; Sellam et al., 2022). However, the effects of randomness influence the choice of hyperparameters as well, possibly leading to suboptimal set of hyperparameters that are determined as optimal only due to random chance. To deal with this problem, some papers search through hyperparameters over multiple runs (similarly to investigation) and choose the best ones based on aggregation over all the runs (Bouthillier et al., 2019, 2021). Using multiple runs to find the hyperparameters guarantees that they are more optimal for the investigation, but at the cost of significant increase in computational cost. However, it is important to note that not all approaches require hyperparameter optimisation (Lu et al., 2022; Winata et al., 2021; Schick and Schütze, 2021; Liu et al., 2022; Zhao et al., 2021a; Efrat et al., 2022; Zhang et al., 2022a, a).
The choice of how many randomness factors to investigate at the same time also plays an important role in the investigation methodology. Many papers consider only a single randomness factor (Reimers and Gurevych, 2017; Bouthillier et al., 2019; Dehghani et al., 2021; Winata et al., 2021; Schick and Schütze, 2021; Liu et al., 2022). However, such investigation can lead to biased results due to the interactions between different randomness factors. Using multiple randomness factors in a single investigation allows for more detailed exploration of the behaviour of models in presence of different sources of randomness. Most straightforward approach used in the papers is performing investigation of single factor multiple times, while only changing the number of samples to observe how the effects change when having more data for training (Agarwal et al., 2021; Setlur et al., 2021a; Ye et al., 2021). However, such investigation considers only the systematic changes introduced by the number of samples and can still be sensitive to the interactions. Most approaches investigate multiple randomness factors at the same time by introducing randomised changes to each one at the same time (Cioba et al., 2022; McCoy et al., 2020; Zhou et al., 2020; Zhao et al., 2021b; Jundi and Lapesa, 2022; Zhong et al., 2021; Efrat et al., 2022; Zhang et al., 2022a; Bouthillier et al., 2021; Boquet et al., 2019; Dodge et al., 2020; Sellam et al., 2022; Lu et al., 2022; Zhao et al., 2021a), while also defining the differences and relations between them (Pham et al., 2021). This allows to take the interactions between randomness factors into consideration to explore their combined effects.
Due to the potentially significant impact of the interactions between randomness factors, it is important to define how to consider these interactions and how to deal with them to disentangle the effects of different randomness factors. Although it is more explicit when investigating multiple factors, it is just as important when investigating only one factor at a time. Many papers do not use any specific way to distinguish between the effects of different randomness factors (Reimers and Gurevych, 2017; Bouthillier et al., 2019; Dehghani et al., 2021; Setlur et al., 2021a; Cioba et al., 2022; Ye et al., 2021; McCoy et al., 2020; Zhou et al., 2020; Agarwal et al., 2021; Winata et al., 2021; Schick and Schütze, 2021; Liu et al., 2022). As mentioned before, this may lead to the results being biased and can be traced as a direct cause of some of the open problems discussed later. A simple way to consider the interactions is to investigate one factor at a time by changing only the configuration of this factor, while using a fixed configuration for all the others (Pham et al., 2021; Zhao et al., 2021b; Lu et al., 2022; Zhao et al., 2021a; Efrat et al., 2022; Boquet et al., 2019). However, such approach is limited in its investigatory strength, due to the randomness still being present in the choice of the fixed configuration. A better way is to search through combination of configurations for all the factors at the same time (Jundi and Lapesa, 2022; Zhong et al., 2021; Zhang et al., 2022a; Bouthillier et al., 2021; Dodge et al., 2020; Sellam et al., 2022). However, this significantly increase the computational cost of the whole investigation (Bouthillier et al., 2021; Dehghani et al., 2021).
Another important aspect to consider is how to obtain representative results from the investigation. The representativeness of results mainly depends how many randomness factor configurations are explored, i.e., how many training and evaluation runs are used during the investigation. Many papers choose the number of different randomness factor configurations only arbitrarily, often choosing lower number (3, 5, 40, 41, 86, 100, 1000) without any explicit explanation why the given number was chosen (Reimers and Gurevych, 2017; Bouthillier et al., 2019; Zhou et al., 2020; Boquet et al., 2019; Ye et al., 2021; Liu et al., 2022; McCoy et al., 2020; Zhao et al., 2021b; Schick and Schütze, 2021; Pham et al., 2021; Efrat et al., 2022; Bouthillier et al., 2021). Similarly, when considering multiple randomness factors and interactions between them, it is common to arbitrarily select some number (usually small) of configurations for each factor and then determine the number of runs as all possible combination of these configurations (Jundi and Lapesa, 2022; Zhong et al., 2021; Dodge et al., 2020; Sellam et al., 2022). For example, when considering 3 different randomness factors, with 5 different configurations for each, the final number of runs is determined as 125 ($5^{3}$). In specific cases, all possible configurations of a single factor can be explored (Dehghani et al., 2021; Agarwal et al., 2021; Lu et al., 2022; Winata et al., 2021; Zhao et al., 2021a; Zhang et al., 2022a). However, this happens only when the number of possible configurations is small or the added computation cost is not as significant.
Majority of the papers evaluate the effects of randomness from multiple runs based on an aggregated value from these runs, either as a single value of mean, standard deviation (with standard deviation not even being present in some papers), or number of failed runs (Bouthillier et al., 2019; Dehghani et al., 2021; Pham et al., 2021; Setlur et al., 2021a; Cioba et al., 2022; Ye et al., 2021; McCoy et al., 2020; Zhou et al., 2020; Zhao et al., 2021b; Jundi and Lapesa, 2022; Zhong et al., 2021; Lu et al., 2022; Winata et al., 2021; Schick and Schütze, 2021; Zhao et al., 2021a; Efrat et al., 2022), or as multiple aggregated values, such as average, worst and best performance, or the difference between them (e.g., difference between worst and best run, or difference between worst and average run) (Agarwal et al., 2021; Liu et al., 2022; Zhang et al., 2022a). Only some papers evaluate the results of investigation based on the distribution, mostly as a means for comparing different randomness factors, models or the behaviour across datasets (Reimers and Gurevych, 2017). Some papers introduce special evaluation strategies based on statistical approaches that can better distinguish between effects of different factors and deal with their interactions, and that can better estimate the results distribution from small number of repeated runs to make the results more representative (Bouthillier et al., 2021; Boquet et al., 2019; Dodge et al., 2020; Sellam et al., 2022).
4.3. Results from the Investigation
Majority of the papers found that the effects of randomness have significant effects on the performance (Bouthillier et al., 2019; Dehghani et al., 2021; Agarwal et al., 2021; Ye et al., 2021; McCoy et al., 2020; Jundi and Lapesa, 2022; Zhong et al., 2021; Lu et al., 2022; Winata et al., 2021; Schick and Schütze, 2021; Zhao et al., 2021a; Efrat et al., 2022; Zhang et al., 2022a; Bouthillier et al., 2021; Boquet et al., 2019; Dodge et al., 2020; Sellam et al., 2022), some specifying how high the variance is, or some simply stating that the variance is significant without providing further details. The significant effects are observed across all approaches for dealing with limited data. Fine-tuning BERT multiple times and picking the best result, it is possible to outperform its more recent variants (Dodge et al., 2020). Changing only order of samples when using language models for in-context learning can lead from state-of-the-art predictions to random guesses (Lu et al., 2022). The difference in performance between worst performing and best performing set of adaptation data in meta-learning can be up to 90% (Agarwal et al., 2021). On the other hand, the observed variance in other experiments was found to be lower, usually only few percent difference (1%-5%), while still being enough to influence the results and comparisons between different models (Pham et al., 2021; Zhao et al., 2021b; Liu et al., 2022; Reimers and Gurevych, 2017).
Almost all investigation is done on in-distribution settings, with only few papers dealing with out-of-distribution evaluation of the stability. When investigating the effects in-distribution, majority of the papers found the variance to be quite high (more than 10%, but sometimes going up to 90% between best and worst performance) (Bouthillier et al., 2019; Dehghani et al., 2021; Agarwal et al., 2021; Jundi and Lapesa, 2022; Zhong et al., 2021; Lu et al., 2022; Winata et al., 2021; Schick and Schütze, 2021; Zhao et al., 2021a; Efrat et al., 2022; Zhang et al., 2022a; Bouthillier et al., 2021; Boquet et al., 2019; Dodge et al., 2020; Sellam et al., 2022), with only few papers finding the variance to be smaller (1%-5% standard deviation) (Pham et al., 2021; Zhao et al., 2021b; Liu et al., 2022; Reimers and Gurevych, 2017). However, all the papers that investigate the out-of-distribution setting find that the variance observed on in-distribution datasets is insignificant (showing  2% variance) compared to the one observed on the out-of-distribution datasets (showing  50% variance in performance (Ye et al., 2021; McCoy et al., 2020; Zhou et al., 2020).
The investigation results indicate how the variance increases or decreases based on choices in the investigation methodology, without specifically addressing the stability. However, there is usually low consensus on how much the variance is reduced by different choices. Some papers found that using more labelled samples helps the stability the most (Agarwal et al., 2021; Jundi and Lapesa, 2022), although with diminishing returns (Zhao et al., 2021b). On the other hand, other papers found that increasing the number of samples does not lead to significant improvement (Lu et al., 2022), or even having negative effects on the stability (Zhao et al., 2021a). In other papers, it was found that keeping the randomness factor configuration fixed during the whole use of the model, such as using a single set of samples or defined order based on heuristics, can help reduce the variance (Setlur et al., 2021a; Winata et al., 2021), but not in all the cases (Zhang et al., 2022a; Boquet et al., 2019). Similarly, some papers found that increasing the model size reduces the variance (Zhong et al., 2021), but others have found that it either has no effect or even worsens the situation (Efrat et al., 2022). Other papers find that accounting for and investigating as many sources of randomness as possible helps with reducing variance at the cost of computational cost (Bouthillier et al., 2021), or that tasks with different difficulty affect the variance in different ways, although it does not hold true that using easier tasks leads to decrease the variance (Cioba et al., 2022; Ye et al., 2021).
The investigation of the effects of randomness is often used just to determine that there is an impact of randomness on the performance and stability of the models, but without providing any additional analysis of how much it affects the models or what randomness factors are most important (Setlur et al., 2021a; Cioba et al., 2022; Ye et al., 2021; McCoy et al., 2020; Zhou et al., 2020; Winata et al., 2021; Schick and Schütze, 2021; Liu et al., 2022; Zhang et al., 2022a; Sellam et al., 2022). It is also observed that randomness can lead to models oscillating between random chance and state-of-the-art results (Lu et al., 2022; Agarwal et al., 2021), or that the randomness can impact the ordering between different models, resulting in simpler models even outperforming more complex ones based only on random chance (Reimers and Gurevych, 2017; Bouthillier et al., 2019; Dehghani et al., 2021; Dodge et al., 2020). The most interesting use for the results of investigation is for comparing between the importance of the effects of different randomness factors. Such investigation results show that the choice of samples and order of data are the most significant contributors to the instability of the models (Efrat et al., 2022; Jundi and Lapesa, 2022; Zhao et al., 2021b), with some results indicating choice being more important (Zhao et al., 2021b; Jundi and Lapesa, 2022) and other that order is the more important (Zhao et al., 2021a). Other results indicate that random seeds, especially the ones used when fine-tuning, are more significant than just the effects of implementation and hardware (Pham et al., 2021; Zhong et al., 2021), while others find the effects implementation and hardware to contribute more significantly (Boquet et al., 2019). Finally, the results from investigation of almost all randomness factors at the same time indicate that the splitting of data has the most impact on the performance (Bouthillier et al., 2021).
The behaviour of randomness factors is compared across different models and datasets. The inconsistency in results is present here as well. Some papers indicate that there is different behaviour of the randomness factors across different model architectures and model sizes (Reimers and Gurevych, 2017; Bouthillier et al., 2019; Dehghani et al., 2021; Zhong et al., 2021; Lu et al., 2022). However, other papers indicate that the behaviour across models is similar, but varies across different datasets (Zhou et al., 2020; Ye et al., 2021; Zhao et al., 2021a), with smaller datasets showing more variance (Dodge et al., 2020). Finally, many of the results from investigation indicate that the behaviour of randomness factors is consistent across all the model architectures, sizes and datasets (Agarwal et al., 2021; McCoy et al., 2020; Zhao et al., 2021b; Jundi and Lapesa, 2022; Winata et al., 2021; Schick and Schütze, 2021; Liu et al., 2022; Efrat et al., 2022; Zhang et al., 2022a; Bouthillier et al., 2021; Boquet et al., 2019; Sellam et al., 2022).
4.4. Overview of Findings: Investigating Effects of Randomness
The effects of randomness are not investigated evenly across all groups of approaches for dealing with limited labelled data or modalities. The investigation is more skewed towards randomness in language models, which goes hand in hand with the focus on natural language processing. However, the findings from meta-learning and overall machine learning approaches indicate that this mostly stems from the popularity of language models research and not from the impact the randomness has on the investigated approaches. The effects of randomness are consistently significant across all approaches for dealing with limited labelled data.
In addition, the focus of the investigation is mostly on a few popular randomness factors, such as random seeds, model initialisation, sample choice and order. Even though other randomness factors receive minimal attention, it has no correlation with their importance. For example, data split is considered the most important factor, while being investigated only in few papers. Another mostly ignored factors are the non-deterministic implementation and hardware, and choice of which samples are labelled.
The relevance and popularity of the randomness factors heavily depend on the investigated approach. Many factors are irrelevant for specific approaches, such as model initialisation for in-context learning. We can also observe different popularity of general randomness factors across different approaches. For example, sample (and task) choice is popular mostly for meta-learning and in-context learning; while investigating the number of samples used is mostly popular in fine-tuning and in-context learning.
The investigation design is agnostic from data and the machine learning approach, making its use consistent across all papers with only small modifications. However, these small modifications have significant impact on the overall investigation, often making it impossible to compare findings from different papers. The number of training and evaluation runs are chosen rather arbitrarily, ranging from ones to thousands, without any explanations behind the choice. Almost no paper introduces heuristics for determining how many repeated runs are needed to estimate the underlying distribution of results. The only heuristic used is to explore all the possible randomness factor configurations when their number is small, such as the permutations when number of used samples is small ($<10$).
A significant drawback for almost all investigation papers is that the effects of randomness are not taken into consideration when searching for hyper-parameters. The papers often do not even mention how the hyper-parameters are chosen, use a fixed set found beforehand in preliminary experiment or in another paper. However, without searching for hyper-parameters over multiple runs, the results may be biased.
The effects of randomness are usually investigated only across small number of randomness factors at one time. The number of considered factors is usually 1-3, while other factors may be mentioned, but not investigated. Almost no paper covers a broader and more representative set of factors at the same time, mostly because it is problematic to investigate the larger set of factors due to the computational cost this introduces - the number of repeated runs raises exponentially with number of randomness factors (Bouthillier et al., 2019, 2021; Dehghani et al., 2021).
Even when evaluating multiple randomness factors at the same time, the interactions between them are not taken into consideration, which leads to biased comparisons. Only some papers consider using fixed randomness factor configuration for the factors that are not investigated at the moment, or using combinations of all randomness factor combinations. However, this is always done only for factors that are considered in the paper and not for all of them. In addition, no paper applies mitigation strategies to other randomness factors.
No analysis of the importance of the randomness factors is performed when investigating a single factor. Similarly, when investigating multiple factors at the same time, no comparison is performed between these factors. This causes the investigation results to be limited in their extent as the problem is only recognised, but no further conclusions can be drawn from it. It is not possible to answer questions like: “Are the effects of the factor significant?” or “Which factor is more important and should be focused on first?”.
The effects of randomness are evaluated in multitude of ways. The use of simple aggregation over the training runs and reporting single standard deviation value is the most common approach. However, as the experiments are performed on a limited number of training runs, the single aggregated value may not be representative of the underlying distribution. A better approach for evaluating the effects is the use of statistical approaches that allow to quantify the uncertainty still present in the models when only limited runs are used, or to determine the expected behaviour of the models in limit, while only having results from significantly smaller number of runs. The evaluation is further complicated by the interactions between randomness factors that can muddle the results and need to be taken into consideration using statistical approaches.
Almost all papers investigate the effects of randomness on in-distribution settings. Even though the papers find significant effects in this setting, the effects found when studying the out-of-distribution settings are more significant. The limited focus on the out-of-distribution settings may be due to limited focus on meta-learning. In addition, in case of fine-tuning and in-context learning, the out-of-distribution settings do not make sense in many cases as the main point is to explicitly calibrate the models to work on the specific task. The only exception is focus on zero-shot learning and cross-lingual approaches.
A significant drawback regarding the results from the investigation is that no obvious consensus exists regarding the behaviour of different randomness factors across different models and datasets. Many papers investigating the same setting come to contradictory findings about almost all parts of the investigation. Some papers observe decrease in variance when the size of the model increases, while others found no obvious connection between size of model and the amount of observed variance (e.g., increasing size lead to increase in variance). When it comes to datasets, some papers observed consistent results, with the same factor being most significant across different datasets, while others observed different factors being important on different datasets. In addition, no external property of the dataset was found that can indicate the amount of variance that would be observed on it. For example, the variance on more complex datasets with more classes and less data was far lower than on simpler datasets with only 2 classes and large amounts of data available.
The most consistent finding from the investigation is that increasing the number of samples reduces variance. However, the decrease in stability has diminishing returns, where from certain point, increasing available data has no more significant impact. This may also be due to the quality of samples used, which was never considered when looking at number of samples used. Although this finding is most consistent, there are also papers that find that increasing number of samples leads to increase in variance in specific settings.
5. Determining the Origin of Randomness
In this task, the origin of randomness is determined, which represents a specific underlying problem that only manifests through the randomness factors. The design of experiments is similar to the one when investigating the effects of randomness, but with a specific aspect that is observed (e.g., whether the label which is used as last example in in-context learning has impact on the prediction, which would indicate recency bias). Analysing the results and behaviour of models on this specific aspect across the different runs, a hypothesis is formed and answered either empirically or using theoretical understanding of deep learning models.
The design of experiments and further explanation of the behaviour is closely tied to the used machine learning approach and the randomness factor that was observed to cause the randomness.
5.1. Origins of Randomness in Meta-Learning
The focus on origins of randomness in meta-learning is limited only to the adaptation data sampling randomness factor (Agarwal et al., 2021). The behaviour of different meta-learning approaches is analysed using combination of all the possible combinations of the adaptation samples from image datasets. The goal of the analysis is to determine the characteristics of the worst case samples, i.e., the samples that cause the worst performance when used for adaptation. It is determined that these samples are not artefacts, as they are always correctly labelled, appear representative of the respective classes and are not visually different from other images. Based on this observation, it is speculated that the problem is due to adversarial nature of the images. However, it is later disproved empirically as adversarial training does not reduce the effects of randomness, leaving the real reason why meta-learning is sensitive to the choice of support samples open.
Another origin of randomness in meta-learning is hypothesised to be overfitting. This can be due to the variance in the training process, specifically in the gradient estimator (Wang et al., 2021a; Yang and Kwok, 2022).
5.2. Origins of Randomness in Language Model Fine-Tuning
Language model fine-tuning is analysed mainly using the random seed randomness factor and in part using choice of samples at the start of training. Many different origins of randomness are identified. The experiments in this part are the most diverse, with the training being performed across different models, languages and even factors, but often leading to similar origins of randomness.
Small availability of data that causes overfitting and catastrophic forgetting is identified as being one origin of randomness when fine-tuning language models (Zhao et al., 2021b; Hua et al., 2021). This can cause the models to rely on shallow lexical features, even though the models were trained on abundant, feature rich data beforehand (Zhao et al., 2021b). Another possible origin may be the bad fine-tuning start point in the transferred model, mainly in the higher layers (not only the newly initialised ones, but also the pre-trained ones) that are sensitive to perturbations in the input.
However, catastrophic forgetting and small size of training data is disputed as cause by later papers (Mosbach et al., 2021; Zhang et al., 2021b; Wu et al., 2022). Instead, it is only a symptom of another problem (Mosbach et al., 2021). Looking at the failed fine-tuning runs (i.e., the runs that perform worse than random chance) it is identified that the language models suffer from vanishing gradients (Mosbach et al., 2021; Zhang et al., 2021b). The use of optimiser without bias correction, further compounded by the use of small number of training iterations, is identified as the cause of vanishing gradients (Mosbach et al., 2021; Zhang et al., 2021b). As opposed to previous papers, it is observed that training for longer does not lead to overfitting, but instead reduces the vanishing gradients problems (Mosbach et al., 2021; Zhang et al., 2021b). Additional cause of vanishing gradients is the sub-optimal initialisation point for the upper layers, which are specialised and cause significant variance when transferred (Zhang et al., 2021b).
Again, all these hypotheses are disputed (Wu et al., 2022) and instead under-specification of the optimisation problem is proposed as the origin of the variance (Wu et al., 2022; D’Amour et al., 2022; McCoy et al., 2020), especially in out-of-distribution datasets (Wu et al., 2022; McCoy et al., 2020). There exists multiple good solutions in the generalisation errors on the source dataset, but with wildly different generalisation errors on the target datasets, with the optimal solution lying in a non-flat region where even a small perturbation leads to large difference in error (Wu et al., 2022; McCoy et al., 2020). This makes the identically trained models to encode different inductive biases (D’Amour et al., 2022) and making the choice of minima rather arbitrary, being easily affected by small changes in initialisation and order of examples (McCoy et al., 2020).
Finally, it is postulated that the large variance in performance is due to the strong inter-example correlations in datasets (Zhou et al., 2020). When the prediction of a single sample changes, this correlation causes simultaneous change in large proportion of the predictions in the dataset. This was observed when analysing the training trajectory in experiment, where the model checkpoints at the same step of training from different initialisation had significant difference in performances. The individual predictions of each sample were highly unstable and changed constantly, which caused the large fluctuations. However, no explanation for the unstable behaviour in individual samples is provided, besides being an inevitable consequence of current datasets (Zhou et al., 2020).
5.3. Origins of Randomness in In-Context Learning
The origins of randomness in in-context learning are analysed across two randomness factors: 1) order of data (Zhao et al., 2021a; Lu et al., 2022); and 2) choice of samples (Liu et al., 2022; Zhang et al., 2022a).
The significant effects of different data orders were found to be the consequence of highly unbalanced label distribution in the output (Zhao et al., 2021a; Lu et al., 2022). The output distribution of the language models is influenced by the biases present in them (Zhao et al., 2021a). Specifically, three biases were identified that cause these effects: 1) majority label bias, where the output is biased towards the more frequent answer in the prompt (e.g., class imbalance in prompt); 2) recency bias, where the model tends to output the labels that appear towards the end of the prompt; and 3) common token bias, where the model is biased to output more common tokens from the pre-training distribution. The combination of these biases causes the model to have highly unbalanced label distribution that needs to be calibrated to reduce the variance (Zhao et al., 2021a). On the other hand, calibrating the distribution in order to reduce the variance and prevent the degenerate behaviour of bad prompts was found to not be effective (Lu et al., 2022). This leads to hypothesis that the biases are not a sole cause of the highly unbalanced label distribution and the origin of randomness should be traced to something different. However, other origins were not found empirically, leading to claims that the sensitivity to ordering of samples is a feature of in-context learning and a core part of every in-context learning approach should be a way to select good performing prompt orderings (Lu et al., 2022).
The random sampling of examples for the in-context learning was determined to be the origin of randomness (Liu et al., 2022; Zhang et al., 2022a) that can also cause problems observed when using different order of examples in prompts. Choosing the samples randomly, without taking similarity or other heuristics into consideration, many low quality demonstrations are selected that may contain spurious correlations and may not provide enough information (Zhang et al., 2022a; Liu et al., 2022). When low quality samples are used, different permutations and calibration of output distribution do not help reduce the variance, which disputes the claims in previous works (Zhang et al., 2022a). Choosing high quality samples based on similarity or heuristics, reduces the variance and allows different ordering and calibration to work (Zhang et al., 2022a; Liu et al., 2022). Therefore, the real origin of randomness, which can also lead to unbalanced distribution and sensitivity to order, is the choice of low quality, dissimilar samples (Zhang et al., 2022a; Liu et al., 2022).
5.4. Overview of Findings: Origins of Randomness
The observed origins of randomness can be summarised as follows:
  • Poor choice of samples that are used for training, adaptation or in-context learning.
  • Overfitting that causes catastrophic forgetting, and models to focus on shallow features.
  • Under-specification, where multiple local minima with same performance are present in training data, which are not consistent with testing data.
  • Highly unbalanced output label distribution stemming from biases present in language models and incorrect ordering of samples.
  • Optimisation problems caused by poor choices in hyperparameters and the problem statement, such as not using bias correction, training for limited number of iterations or poor initialisation and re-use of some neural network layers.
The individual origins of randomness are closely tied to the machine learning approach and the randomness factor investigated. This is mainly the consequence of the experiments that are designed to observe the behaviour of different approaches. Although many different origins are identified, all are tied to small number of randomness factors, specifically the choice of samples, order of the samples in training and the model initialisation.
The popularity of different machine learning approaches and randomness factors is clearly shown in this task for addressing the effects of randomness. The focus on determining the origins of randomness in meta-learning is limited, while language model fine-tuning is most popular.
The first effects of interactions between randomness factors are observed. In many cases, the different origins of randomness identified are later disputed and denoted as a simple consequence of other effects of randomness. For example, it is observed that choice of samples in in-context learning causes problems that are often incorrectly attributed to the ordering of these samples.
This further shows the prevailing inconsistency of results from different papers. The compounding effects between different randomness factors cause significant problems when not properly addressed. Another aspect that plays role is the simple evaluation, as the interactions are clearly shown when more comprehensive evaluation is utilised.
6. Mitigating the Randomness
In the mitigation task, the effects of randomness are mitigated to reduce the variance as much as possible and to improve the overall performance using mitigation strategies. We divide these mitigation strategies into two separate groups: 1) general mitigation strategies; and 2) problem-specific mitigation strategies.
The general mitigation strategies can be used to mitigate the effects of randomness stemming from any origin of randomness and randomness factor. So far, the only way to deal with any origin of randomness is to repeat the training process multiple times, each time with different randomness factor configuration (called multi-run mitigation strategy). Even though such strategy do not really mitigate the variance, the idea is that aggregating over enough repeats of the training runs leads to more objective results, as the underlying distribution can be better estimated and the expected behaviour can be better derived from it (Bouthillier et al., 2021; Dehghani et al., 2021).
Although the multi-run mitigation strategy is implicitly used in investigation and recognise papers (due to the repeated runs), it is explicitly mostly used in combination with the data split randomness factor, as it can be viewed as a modification of the cross-validation. Using this strategy, multiple random data splits are macro-averaged together to produce more objective results (Mukherjee et al., 2021). The strategy is also modified to increase its power and reduce the number of repeated runs needed. One such modification is the use of fixed ratio splitting without constraints instead of cross-validation (Zheng et al., 2022). The fixed ration splitting can better approximate unseen data as some samples appear only in training or test split of data. Even though it is mostly used with the data split randomness factor, it can be extended for use on any randomness factor. In some cases, it is the only available solution for mitigating the randomness, i.e., when the randomness factor cannot be easily controlled, such as the effects of implementation of deep learning frameworks and hardware.
On the other hand, the problem-specific mitigation strategies are designed to deal with a specific origin of randomness, making them dependent on the randomness factor and the origins of randomness as analysed in Section 5.
6.1. Poor Choice of Samples
The variance in performance resulting from the choice of samples can be introduced at 2 separate levels of the training process: 1) at the start of training, when sampling a sub-set of representative samples that are used during training; and 2) during a single step of training, when sampling the data used for the specific training iteration. The first level represents the real world scenario where there is only a limited labelling budget available that needs to be effectively utilised by choosing samples with as much information as possible. By definition, the choice at the start of training process should have a significant influence on the choices in the individual training iteration, leading to interactions between these two levels. In case of fine-tuning, both choices can be merged into one, as all the training samples are usually used at each step of training.
6.1.1. Selecting Representative Samples
When choosing representative samples for training, each proposed approach assumes the existence of a large set of unlabelled data from which the samples are chosen for annotation (Chang et al., 2021; Zhang et al., 2022a). When selecting only small number of samples ($<100$), a simple clustering can be used to select the representative samples (Chang et al., 2021). To choose K samples, the whole set of unlabelled data is clustered using K-means clustering into K clusters. The samples closest to the centre of each cluster are selected to form the set of K samples for training (Chang et al., 2021). To maximise the probability that the clusters are spread evenly across the data, the initial centroids are chosen based on the probability proportional to the squared distance from all the other existing centres instead of choosing them randomly (Chang et al., 2021). Another possibility to select the samples is to use reinforcement learning, similarly to active learning (Zhang et al., 2022a).
The problem of effective sample choice is also present in cross-lingual learning, when choosing what samples to translate to different languages. One proposed solution is to train a linear model that determines the gain in performance and stability, when a sample with given characteristics is translated and used for training (Jundi and Lapesa, 2022). As dataset for training this linear model, the performance scores from a large set of cross-lingual experiments where random samples are manually translated is used. The linear model is trained to predict the performance scores using language features and similarity scores between embeddings (Jundi and Lapesa, 2022). Afterwards, the model can be used to select which samples to select for translation based on the added gain (Jundi and Lapesa, 2022).
6.1.2. Sampling Data for Training Iteration
To select better samples for training iteration, the proposed solutions use heuristics instead of random sampling (Liu et al., 2022; Agarwal et al., 2021). One possibility is to select samples based on a similarity using k-nearest neighbour (Liu et al., 2022). Selecting samples closest to the test ones can significantly increase the stability without incurring significant computation overhead when number of samples is small (Liu et al., 2022). Another possibility is to use adversarial training on the worst-case adaptation samples (found using greedy search) in meta-learning (Agarwal et al., 2021). However, it provides no significant improvement in regards of stability for specific meta-learning approaches (Agarwal et al., 2021).
Last, but not least, curriculum learning can be used to select optimal samples for each training step (Zhang et al., 2021a). In this approach, the difficulty of samples is first determined using several teacher models that are trained independently on the data. These teacher models separate the data into multiple buckets based on their difficulty. Afterwards, the instances are sampled from the buckets, starting from the easiest and gradually increasing in difficulty.
6.2. Overfitting and Catastrophic Forgetting
How the overfitting and catastrophic forgetting is mitigated is closely tied with the approach for which it is designed. Therefore, we provide further categorisation based on these approaches
6.2.1. Overfitting in Meta-Learning
One solution to reduce the overfitting in meta-learning approaches is the augmentation of the support and query data, and the task construction (Osahor and Nasrabadi, 2022). Applying the augmentation to the support and query data increases the number of samples for training which improves the performance by making the training more robust. The augmentation applied to the tasks increases the number of classes per individual tasks trained on, improving the performance as well by providing more diverse tasks and improving generalisation (Osahor and Nasrabadi, 2022).
Another solution is the use of variance reduction algorithms in the optimisation based meta-learning (Wang et al., 2021a; Yang and Kwok, 2022). One possibility is to focus on first-order meta-learning algorithms and introduce the variance reduction only to them (Wang et al., 2021a). The variance reduction term is introduced to the gradient estimator in task adaptation stage, motivated by the recursive momentum technique in (Cutkosky and Orabona, 2019). This term is initialised by averaging the gradients from randomly sampled tasks using the initial parameters and then is further updated using a weighted sum of mini-batch stochastic gradients across all sampled tasks (Wang et al., 2021a). As this foregoes the benefits of the bi-level meta-learning process, other papers propose to modify the variance reduction term (STORM (Cutkosky and Orabona, 2019; Levy et al., 2021)) for use on bi-level meta-learning optimisation, in combination with large number of training steps to reduce the amount of storage required (Yang and Kwok, 2022). To achieve this, the variance-reduced gradient estimate is computed using stochastic gradients from two successive iterates, while evaluating the gradient using the previous two iterates on the current batch of samples (Yang and Kwok, 2022).
6.2.2. Overfitting in Language Model Fine-Tuning
The augmentation technique, in combination with regularisation, is also proposed when fine-tuning language models (Meng et al., 2022). The pre-trained language model is used as a generator for new data samples used for augmentation. To guarantee that samples with discriminative labels are generated, a meta-weighted maximum likelihood objective can be used for the tuning of the generator. As the resulting labels still contain some level of noise, a noise-robust procedure for training the final model is proposed as regularisation using label smoothing and temporal ensembling (Meng et al., 2022).
Another solution is to look at the disconnection of problem statement between pre-training and fine-tuning stage. It is common occurrence that the objective and even the input and output distributions are different which causes the fine-tuning process to be brittle, especially when only small amount of data is available (Phang et al., 2019). A second pre-training step on a data-rich supervised tasks can be used to better prepare the model for fine-tuning stage, mitigating the brittleness and tendency to overfit on data (Phang et al., 2019).
Last set of approaches utilises regularisation to prevent overfitting. Injecting Gaussian like noise vector into the input can be used as a stability regularisation, similar to augmentation (Hua et al., 2021). Another regularisation technique is to perform fine-tuning update only on a sub-network instead of the whole network (Zhang et al., 2018; Xu et al., 2021c; Zhang et al., 2022b). Mixup, introduced in (Zhang et al., 2018), randomly selects sub-network with a given probability and replaces the parameters with pre-trained weights, while updating the rest of the network. Instead of random selection, CHILD-TUNING, introduced in (Xu et al., 2021c), uses Fisher Information to determine the importance of parameters based on all samples before training. Afterwards an unchanged sub-network with the most important parameters is selected for the update. The previous two approaches can be improved using Dynamic Parameter Selection, introduced in (Zhang et al., 2022b), which adaptively selects a promising sub-network to perform the fine-tuning steps on. The sub-network is chosen dynamically, i.e., different sub-network is potentially fine-tuned at each training step, based on importance of the parameters. It greatly reduced the heavy computation costs introduced by CHILD-TUNING (Xu et al., 2021c) that separate the process of fine-tuning and decision which sub-network to optimise. In addition, it can better capture the task-relevant parameters across sub-networks of different sizes and bring stable improvement in various parts of the model.
6.3. Under-Specification
To deal with under-specification, models with stronger inductive biases that can help distinguish between many local minima should be proposed (McCoy et al., 2020; Zhou et al., 2020), or the training sets should be better designed (McCoy et al., 2020; Zhou et al., 2020). The better sets should represent the phenomena appearing in texts to reduce the probability of local minima not transferring to testing data (McCoy et al., 2020). In addition, the tasks used for large scale pre-training should adhere to structure and level of compositionality (i.e., solving combination of multiple task leading to the solution for a different, more complex task), and more diverse datasets in terms of syntax and lexicon should be constructed (Zhou et al., 2020).
Another solution for dealing with under-specification is to utilise an ensemble of models (Hidey et al., 2022; Khurana et al., 2021). One possibility to construct the ensemble is to combine predictions from multiple trained models using their uniform average (Hidey et al., 2022). Another possibility is to use stochastic weight averaging (SWA) as regularisation for the ensembles (Khurana et al., 2021). Instead of training the model multiple times, SWA creates the ensemble by averaging over different snapshots of model weights. Such ensemble explores other solutions around the local minima, while also reducing computation cost as the ensembling at weight level needs only a single training run (Khurana et al., 2021). Although the SWA method consistently achieves better stability of performance than standard fine-tuning when used over different values of random seeds, the increase in stability is only slight (Khurana et al., 2021). In addition to ensembling, distillation and co-distillation can also be used (Hidey et al., 2022). Distillation can be viewed as the exact opposite of ensembling, where the knowledge of an ensemble of models (or a single model) is compressed to less number of parameters by training a student model on the prediction of a pre-trained teacher model. In co-distillation, multiple peer models are trained in parallel instead, each sharing architecture and training data. Each of these techniques reduces the variance in results, improving stability (Hidey et al., 2022).
6.4. Highly Unbalanced Label Distribution
To deal with highly unbalanced label distribution in the in-context learning it is proposed to use calibration (Zhao et al., 2021a; Min et al., 2022). One possibility is to use contextual calibration, where the output distribution is first estimated using neutral prompts (Zhao et al., 2021a). The neutral prompts contain only in-context samples with the testing sample having its text replaced with “N/A”. Using multiple such prompts, the distribution between labels can be estimated and then corrected using affine transformation (Zhao et al., 2021a). For example, if the model outputs positive labels in more than 50% of cases (e.g., in 65%) in binary classification, the testing example is considered to be positive only when the confidence is higher than this threshold (e.g., more than 65%). Instead of using neutral prompts, it is possible to parameterise the distribution using Bayes rule (Min et al., 2022). To use it, the language model is not presented with K concatenated samples in a single prompt, but instead presented one sample at the time. Obtaining output probabilities and multiplying them together can be used to estimate the distribution, while also reducing the dependency on ordering of samples (Min et al., 2022). Compared to direct distribution estimation, this provides larger improvements in terms of stability when there are fewer training samples, the data is imbalanced, the number of classes is large, or a generalisation to unseen labels is required (Min et al., 2022).
Apart from calibration, the highly unbalanced distribution is addressed by finding the best ordering of examples (Lu et al., 2022; Kumar and Talukdar, 2021). One possible approach to automatically generate good prompt orderings is to use heuristics and entropy (Lu et al., 2022). A probing set is first constructed from all permutations of randomly selected set of training samples that are used to query the language model. Each ordering in the probing set is scored using local and global entropy without the need for labelled test data. This produces ranking for all the orderings, allowing the choice of best order of samples that avoid the issue of unbalanced predictions (Lu et al., 2022). Another possibility is to search over possible permutations of training samples to find the best performing one using genetic algorithm (Kumar and Talukdar, 2021).
6.5. Optimisation Problems
The mitigation strategies for dealing optimisation problems in language model fine-tuning are simply a specific fixes for the poor choices that introduce the effects of randomness. One of the fixes is to increase the number of iterations considerably and introduce a bias correction and small learning rate to avoid vanishing gradients (Mosbach et al., 2021; Zhang et al., 2021b).
Another solution is to modify the way the additional layers are initialised (Dauphin and Schoenholz, 2019). Instead of initialising the layers randomly, which can lead to bad initialisation, a meta-learning approach can be used to find an initialisation that is amenable to learning using gradient descent (Dauphin and Schoenholz, 2019). In addition, badly performing initialisation can be detected as early as the second training iterations thanks to the high correlation of performance at the start and the end of the fine-tuning process (Dodge et al., 2020). Starting many runs with different initialisation and stopping the ones with bad performance in the early stages of fine-tuning, it is possible to search through the whole space of random initialisation to find the best performing ones without any significant increase in computation cost (Dodge et al., 2020).
Yet another solution is to also randomly re-initialise top layers of the language models, in addition to the classification layer (Zhang et al., 2021b). The number of re-initialised layers depends on the problem statement and the model, with more complex models and datasets benefiting from larger number of re-initialised layers, but helps only up to a certain point (Zhang et al., 2021b).
To better deal with the sensitivity to sample choices, ordering of samples and to deal with biases, the problem of in-context learning can be reformulated as a meta-learning problem, i.e., in-context learning is combined with optimisation meta-learning (Chen et al., 2022). The language model is directly optimised to perform the simple sequence prediction on variety of tasks, learning to better perform well on all of them. This allows the model to perform better at test time, while requiring only prompts with instruction and few samples without any further adaptation (Chen et al., 2022).
6.6. Overview of Findings: Mitigating Origins of Randomness
The proposed mitigation strategies are closely tied to the origin of the randomness. As they are explicitly designed to deal with the specific origin of randomness and the randomness factor it was observed around, the proposed strategies are unusable for other randomness factors and origins of randomness. The multi-run mitigation strategy is the only one usable on any randomness factor. Due to this, some randomness factors can be mitigated only using the multi-run strategy, such as the non-deterministic implementation of deep learning frameworks and hardware.
Even though the mitigation strategies address specific origin of randomness, many papers do not provide any kind of extensive analysis of where the randomness originates from. In some cases, the works reference other papers that found this origin, but in many cases no explanation is provided why the mitigation strategy targets the specific origin of randomness.
Each mitigation strategy incurs additional computation cost. The most significant increase is introduced by the multi-run mitigation strategy as it requires large number of model training and evaluation runs. The expected computation cost is several times higher than in mitigation strategies that require only single training run. So far, no general mitigation strategy exists that could provide mitigation without high computational costs. In case of problem-specific mitigation strategies, the computation cost depends on multiple aspects, such as how often it is applied (once before training run or at each training step), how expensive is the approach used (genetic algorithm or simple clustering), or how many repeated runs are needed (e.g., when ensembling is used).
Evaluation of the effectiveness of the mitigation strategies is done in a simple fashion by comparing the performance metric of the model after applying the mitigation strategy. The comparison is done on a single aggregated value, often with low number of repeated runs (sometimes even one), while foregoing the reporting of standard deviation or confidence intervals. In addition, the interactions between factors are completely disregarded.
Such simple evaluation that disregards the interactions may be the main reason for inconsistency in results. Similar to the investigate and determine tasks, multiple papers report contradictory findings. The mitigation strategy that works in one paper provides no benefits in other.
Finally, only a single mitigation strategy is applied at one time in most papers. The papers that use multiple mitigation strategies mostly use heuristics to deal with optimisation problems, such as increasing the number of training runs and introducing bias correction into the optimiser.
7. Comparing and Benchmarking: The Impact of Randomness in Low Data Regimes
In this task, the effects of randomness are taken into consideration when comparing and benchmarking different approaches for learning with limited labelled data. As the effects of randomness can have significant effect when comparing different approaches, or when designing benchmarks, modifications to the process of benchmarking and comparing is required. Even though the benchmarks and comparisons between approaches are closely related, the proposed modifications that can deal with effects of randomness look at the problem from different perspectives.
When simply comparing the approaches, the modifications are introduced at the end of training. The design of the modified comparisons is independent from the training process and, in theory, can be applied on any dataset, model, or even in benchmarks. The independence from the training process makes the modifications easier to design, but often require specific assumptions, such as performing multiple runs. The modifications are usually designed to be simple, often taking form of best practices for the comparisons (e.g., use statistical significance tests on distribution of results). An important aspect for the comparisons is what metrics are reported.
On the other hand, the benchmarks are specifically designed to take the randomness into consideration from the start of training. The whole training process is modified based on desiderata that address the randomness, making benchmarks closely tied to the training process. Even though it makes the benchmarks harder to design, they often provide better explanation and comparative power as no assumption and constraints are needed.
7.1. Comparing Approaches in Low Data Regimes
Comparisons between different approaches in low data regimes are significantly affected by the effects of randomness. Comparing on single value, the ranking of models can vary significantly across repeated training runs that differ only in the randomness factor configurations (Dehghani et al., 2021; Dodge et al., 2020; Reimers and Gurevych, 2017). However, using multiple runs and comparing based on an aggregated values from these runs, with or without standard deviation, is also identified as non-satisfactory, as it cannot easily deal with the variance (Zhong et al., 2021; Dror et al., 2019). If a model achieves better aggregated performance, but with higher standard deviation, it cannot be confidently concluded that it is significantly better.
To deal with the effects of randomness, multiple best practices and modifications to evaluation practices are proposed. All of these modifications and best practices propose to compare on multiple evaluation runs to achieve unbiased comparisons (Schick and Schütze, 2021; Reimers and Gurevych, 2017; Bouthillier et al., 2019; Dehghani et al., 2021; Sellam et al., 2022; Bouthillier et al., 2021; Dror et al., 2019).
A simple modification is to select a set of samples beforehand for each run and compare the models on the same set of fixed samples (Schick and Schütze, 2021). Although it reduces the bias in results, the comparison is still done on a single aggregated value. Therefore it may not be possible to confidently draw conclusions about the difference in the behaviour of the models.
To capture the behaviour of different models, the comparison should be done based on their score distribution from multiple runs instead of a single aggregated value (Reimers and Gurevych, 2017; Bouthillier et al., 2019; Dehghani et al., 2021; Sellam et al., 2022; Bouthillier et al., 2021; Dror et al., 2019). Treating the results from multiple runs as random samples, it is possible to compare the models using statistical significance testing (Dehghani et al., 2021). Drawing conclusions from such comparisons reduces the risk of rejecting promising approaches or falsely accepting weaker ones (Reimers and Gurevych, 2017). In addition, comparing score distributions from multiple runs allows to identify situations in which the models fail and to identify the shared properties between different models on the corresponding tasks that can be used for further comparisons (Bouthillier et al., 2019; Zhong et al., 2021). For example, architectures that differ only in number of parameters can be compared based on the fraction of decaying instance (i.e., instances on which models with larger number of parameters are worse) between models of different sizes (Zhong et al., 2021).
However, comparing between score distributions may not be as straightforward in all cases (Bouthillier et al., 2021; Sellam et al., 2022; Dror et al., 2019). To provide truly unbiased comparisons, as many sources of randomness (i.e., randomness factors) as possible should be randomised, while exploring different sets of hyper-parameters for the models (Reimers and Gurevych, 2017; Bouthillier et al., 2021). This not only leads to better evaluation of the variance associated with each source of randomness, but the error of the expected performance is also reduced. However, such comparisons start to become problematic when the evaluated models are not under our control (i.e., behind API) and the only controllable variance is the data we use. In addition, accounting for all sources of randomness is often not feasible due to the computational cost of running training multiple times (Bouthillier et al., 2021). A statistical bootstrap approach can be used to estimate the distribution of results using repeated sampling of the results, without the need for extensive training runs (Sellam et al., 2022). Finally, even statistical tests may not compare the models with sufficient confidence (as only a sample of score distribution is observed) when not adhering to more strict set of criteria (Dror et al., 2019).
Besides simple metrics, multiple papers propose to report additional ones that can allow for more unbiased comparison using statistical tests. Such metrics include standard deviation, maximum, minimum, mean and confidence intervals (Schick and Schütze, 2021; Bragg et al., 2021), the score distributions and individual episode results from multiple runs, data splits, or even across different datasets (Dehghani et al., 2021; Dror et al., 2019; Zheng et al., 2022; Mukherjee et al., 2021; Bragg et al., 2021), or the computation time and memory required (Bouthillier et al., 2021). This also includes assumptions and constraints of the models, such as reporting whether the model is designed for out-of-distribution evaluation and should not be used in-distribution (Setlur et al., 2021b). Last, but not least, reporting situations in which models fail and properties of models on different tasks should help with comparing the models based on the implicit biases and in identifying parts that can still be improved (Bouthillier et al., 2019).
7.2. Benchmarks Taking Randomness Into Consideration
Multiple benchmarks designed for evaluating few-shot learning are starting to take the problem of randomness and its effects on the stability into consideration (Bragg et al., 2021; Mukherjee et al., 2021; Setlur et al., 2021b; Ye et al., 2021; Zheng et al., 2022; Efrat et al., 2022). However, the extent to which the randomness is addressed is different across the benchmarks.
In the most simple case, only some of the properties related to stability and/or randomness are included in the design of the benchmark (Bragg et al., 2021; Mukherjee et al., 2021; Setlur et al., 2021b; Ye et al., 2021). This includes focus on variable number of shots and classes in few-shot learning to better simulate real world scenarios and to test robustness of models to these systematic changes (Bragg et al., 2021), comparing the performance from multiple runs with the human performance (Mukherjee et al., 2021), reporting of additional properties such as confidence intervals, standard deviation and individual results from different episodes of training to allow for paired statistical tests (Bragg et al., 2021), providing means to evaluate both in-distribution, as well as out-of-distribution (Setlur et al., 2021b), or looking at the behaviour of models in cross-task settings (Ye et al., 2021).
Later benchmarks also explicitly encourage the evaluation on multiple training runs (Mukherjee et al., 2021; Ye et al., 2021; Zheng et al., 2022; Xu et al., 2021b). Such benchmarks mostly provide multiple splits of data for training (Mukherjee et al., 2021; Ye et al., 2021; Zheng et al., 2022; Xu et al., 2021b), while providing either a single set of test data (Mukherjee et al., 2021; Xu et al., 2021b) or providing multiple test data splits (Ye et al., 2021; Zheng et al., 2022). To explore the behaviour in cross-task settings, the splitting is also done on the level of tasks, with some tasks being used only for testing purposes (Ye et al., 2021).
Yet further benchmarks focus on accurately estimating the differences between approaches by taking into consideration the effect of randomness on hyper-parameters (Zheng et al., 2022). Instead of using a fixed set of hyper-parameters, which was found to be sub-optimal, a search across the hyper-parameter space is done over multiple splits of data. To allow for better search over the hyper-parameter space, the splitting strategy is modified based on three desiderata (Zheng et al., 2022): 1) a good data split strategy should result in optimal hyper-parameters on test set; 2) development and test sets should be correlated over the whole hyper-parameter space distribution, i.e., changing hyper-parameters should have no effect on performance correlation between development and test dataset; and 3) changing number of splits should have no impact on the performance and its correlation.
Finally, some benchmarks are specifically designed to determine the sensitivity of the approaches to the effects of small, random changes (Efrat et al., 2022). Instead of focusing on large and complex tasks, the benchmark introduces a compact set of tasks that are short, can be quickly and automatically evaluated, and are trivial for humans, i.e., tasks that elementary school students should perform perfectly at. This includes tasks such as writing a sentence containing a specific word (which is mentioned in instructions), or choosing which of the 2 or more words are longer. The effects of randomness are directly estimated by changing the order of specific words in instructions (e.g., switching the words whose length should be compared), replacing some words (e.g., generate sentence containing word ”cat” instead of ”dog”), changing instruction template, or measuring the gap between two similar tasks. To further avoid confounding effects in stability, the evaluation is performed in zero-shot setting. By multiplying the achieved accuracy with the mean sensitivity score (which is calculated across the different small, random changes), the benchmark explicitly penalises the models for their brittleness.
7.3. Overview of Findings: Comparisons and Benchmarks
The comparisons are specifically designed to deal with the significant effects of randomness when dealing with limited labels. Even when comparing across multiple training runs, it was observed that using a single aggregated value leads to biased comparisons. The most common modification is to compare between result distributions using statistical tests instead. Although this leads to more unbiased comparisons, there are still problems present.
Such comparisons are sensitive to the number of training runs, as the distribution of results can be objectively determined only with larger number of runs. However, running multiple training runs was found to be infeasible in many papers. To deal with this problem, approaches that approximate the distribution without the need for many training runs were proposed recently, such as statistical bootstrapping.
The modified comparisons are utilised only for comparing between different approaches across datasets and tasks. However, it is important to also compare between effects of randomness factors, while taking the interactions into consideration. Even though not specifically designed for it, the modifications should be applicable to this problem as well.
When it comes to benchmarks, the randomness is only starting to be taken into consideration, leading to very simple ways to address it. Most benchmarks simply evaluate across multiple splits or training runs, while keeping the number of runs quite small, or simply explore some properties related to stability. In addition, even though multiple training runs are performed, many benchmarks compare the models only on a single aggregated value. Only a single benchmark explicitly measures the sensitivity of models to effects of randomness and uses it to penalise the final performance metric.
8. Challenges, Open Problems and Next Steps
Based on the findings from the comprehensive literature overview presented in previous chapters, we identify seven challenges and open problems for addressing the effects of randomness on stability of learning with limited labelled data.
8.1. Focus on Small Fragment of Randomness
The scope of investigating the effects randomness, determining its origin and the mitigation of these effects is limited in its extent. The problem can be observed on multiple levels:
  • Approaches and modality - the focus is mostly on the approaches utilising language models (fine-tuning and in-context learning), while focus on other machine learning approaches, such as meta-learning, is severely limited in its extent. The majority of the focus is on natural language processing, as it is the modality used in language models, limiting the analysis of behaviour on other modalities. Therefore the effects of randomness are addressed only on a specific sub-part of learning with limited labelled data, even though the randomness also has significant effects on other machine learning approaches (investigated in this survey).
  • Randomness factors - many factors receive limited attention, such as the effects from splitting data, choosing what data is labelled, or the impact of non-deterministic implementation and hardware. Instead, the focus is on factors in later stages of the training process, such as ordering the samples and model initialisation (hidden behind random seed). This limits the analysis of randomness to a few popular randomness factors. The under-researched factors may still be important, but there is no consensus due to the limited focus. Some papers claim the effects of these factors to be insignificant (Bouthillier et al., 2019), while other consider them to be the most significant contributors of variance (Bouthillier et al., 2021; Zheng et al., 2022; Mukherjee et al., 2021).
  • Number of randomness factors - the experiments are designed to jointly analyse only a small number of factors (1-3). Such experiments compound the limited focus on some factors, while negatively impacting the possible findings from experiments, such as determining the importance of different factors or their interactions. In addition, focusing only on specific, low number of factors leaves the results of the investigation open to the still present effects of other sources of randomness that may skew the results.
We argue that the scope should be widened across all tasks for addressing the effects of randomness, to better explore the behaviour of different machine learning approaches across different modalities and randomness factors. A simple extension would be to perform multiple experiments, each for different randomness factor across all the different approaches the factor is relevant for. A more in-depth analysis would be to analyse the effects of randomness across all factors in single experiment. However, increasing number of randomness factors that are considered in the investigation significantly increase the computation cost (Bouthillier et al., 2021). To overcome this problems, more sophisticated experiments should be designed.
8.2. Limited Analysis and Explanation of Importance for the Effects of Randomness Factors
A significant part missing when investigating the effects of randomness is a thorough analysis of the importance for the different randomness factors. In many cases, the importance of randomness factor is obvious, such as when it causes large variance in results ($>10\%$). When the observed variance is smaller, an analysis and an explanation why the difference in performance is significant and why the effects of the randomness factor are important is required. However, such analysis is mostly missing in the investigation experiments. The investigation should provide us with a level of understanding for the importance of the randomness factors. Especially when determining the effects as single aggregated value. The effects measured as 1% standard deviation have different significance when the performance is in high 90% than when it is in low 20%. However, almost all papers simply present a single aggregated value without further details.
In addition, it is common to work with limited computational budget in real world scenarios (e.g., as a a part growing effort of green AI that aims to limit the CO2 footprint of the AI). Analysing different randomness factors and comparing the importance of their effects would allow us to identify which factors are more important. Afterwards, we can identify how best to allocate our resources to dealing with them, such as ignoring the less important ones. Only few papers compare between factors, such as (Pham et al., 2021; Zhao et al., 2021b; Jundi and Lapesa, 2022; Zhong et al., 2021; Zhao et al., 2021a; Bouthillier et al., 2021, 2019).
Furthermore, the importance of randomness factors may not be consistent across different approaches and modalities. However, there is no consistent information whether this is the case, as the comparisons between effects of randomness factors are limited, or even non-existent.
8.3. Disregarding Interactions Between Randomness Factors
Ignoring interactions between different randomness factors can lead to biased results. We can observe that many papers often find contradictory findings across all the tasks for addressing the effects of randomness. The effects of interactions are most notable when evaluating the mitigation strategies, as this task has the least developed evaluation strategies, often using simple evaluation on a single randomness factor. We argue that the interactions between randomness factors are one of the primary contributors to the inconsistency in findings.
The interactions should be taken into consideration in every task for addressing the effects of randomness, whether it is investigation of the effects, identifying the origin of randomness, proposing mitigation strategies, or even comparing and benchmarking different randomness factors and areas of machine learning.
One possibility is to include all randomness factors in the process, such as exploring through all the combinations of the randomness factor configurations for each randomness factor. However, this incurs significant computational cost due to the number of runs growing exponentially with the number of randomness factors and so may be infeasible. Some papers already do this, but only for a set of selected randomness factors with low number of randomness factor configurations (Zhao et al., 2021a; Jundi and Lapesa, 2022; Zhao et al., 2021b).
Another possibility is to mitigate the effects of randomness factors that are not currently investigated using mitigation strategies. However, such approach has not been utilised so far.
Taking interactions into consideration also has impact on the evaluation of the effects of randomness, estimation of the mitigation strategy effectiveness and comparisons between approaches and randomness factors. Each such tasks need to be explicitly modified to take the interactions into consideration. Specifically, the effects of different randomness factors need to be disentangled. Statistical approaches, designed in some papers, that can disentangle these effects already show promising results for dealing with this problem (Bouthillier et al., 2021; Dodge et al., 2020).
8.4. Effects of Randomness Evaluated Using Basic Techniques and Metrics
Majority of the papers evaluate the experiments when addressing the effects of randomness in a basic manner. Besides evaluating using a single aggregated value from multiple runs, the number of runs are chosen arbitrarily, which causes skewed estimation of the real distribution of results. Therefore, the evaluation is open to the still present effect of randomness which can lead to biased results. Again, such effects can be the source of the inconsistency in results across different papers.
More sophisticated evaluation strategies (based on statistical approaches) that can account for the still present uncertainty in the limited runs, or the ones that can better estimate the real distribution of results without need of large number of samples (training runs), need to be designed to deal with this problem. Currently, only few works determine the effects of randomness based on such statistical approaches (Bouthillier et al., 2021; Dodge et al., 2020; Sellam et al., 2022). At the same time, these evaluation strategies need to be used throughout all the tasks for addressing the effects of randomness. Even though some more sophisticated approaches already exist, many tasks (such as evaluating the effectiveness of mitigation strategies) still use a basic evaluation based on aggregated value, often from a single run.
8.5. Inconsistency in Findings from Different Tasks for Dealing with Effects of Randomness
As already indicated, there are many contradictory findings across papers. In some cases, it is observed that the randomness factors behave in similar fashion across datasets, modalities and architectures, while in other cases their behaviour is inconsistent. We already mentioned two possible sources of this inconsistency, specifically the interactions between randomness factors that are not taken into consideration and simple evaluation across low number of runs. As another contributor we identify the inconsistent investigation methodology.
Even though the effects of randomness are mostly investigated using a consistent approach (i.e., running multiple training and evaluation runs), there are aspects of the investigation methodology that differ across papers. This includes aspects like how to handle hyper-parameters, how to design the investigation experiments, how many runs to perform, or how to evaluate the results. Using different choices for these aspects prevents comparison between results from different papers. It is analogous to benchmarking, where the specifics of how we train the models (e.g., choosing model hyper-parameters, or the number of training iterations used) play a significant role.
The different important aspects should be identified and made as consistent as possible, or be explicitly described to deal with this problem when investigating the effects of randomness. A general methodology, or at least its template, should be created. Such methodology should, among other things define:
  1. (1)
    How to handle hyper-parameters? Should the hyper-parameters be found beforehand for each model and then used for the whole process, or should we search for optimal hyper-parameters at the start of each repeated run? In addition, we should answer the question how to find the optimal hyper-parameters while taking the effects of randomness into consideration. A search over a single set of fixed randomness factor configurations may not be the best idea, but searching over repeated runs introduce significant computational overhead that we may not afford.
  2. (2)
    How many repeated runs to run? Choosing the number of repeated runs will result in better comparability between different investigation runs. However, arbitrarily choosing a low number of runs and then sticking to it across all investigations is not an optimal solution, as it can lead to biased results due to non-representative distribution of the results.
  3. (3)
    How to design the investigation strategy? It should specify how to handle the interactions between randomness factors. For example, if we choose to mitigate other factors it should specify what mitigation strategy is used.
  4. (4)
    What evaluation method to use? Multiple evaluation methods exist, each leading to potentially different findings based on how they handle interactions and multiple runs.
Following the analogy of benchmarks, we imagine that many different investigation frameworks can be proposed, with their designs reflecting some specific aspect deemed to be important (e.g., handling interactions, comparing effects of different randomness factors, computational cost). However, so far no such explicit, fully fleshed out framework exists, only starting points of ones, mainly in papers introducing better evaluation strategies (Boquet et al., 2019; Bouthillier et al., 2021; Dodge et al., 2020; Sellam et al., 2022).
8.6. Absence of Effective Mitigation Strategies for Multiple Randomness Factors
Many partial solutions for mitigating the effects of different randomness factors exist. Such mitigation strategies are explicitly designed to deal with specific origin of randomness and are always used one at a time. However, a strategy that can address all randomness factors, along with the interactions between these factors effectively, is still lacking. In addition, many randomness factors have no specific mitigation strategy designed for them. In some cases, a specific mitigation strategy may not even be possible, such as when dealing with non-deterministic implementation and hardware.
The only mitigation strategy that is usable for all randomness factors is the multi-run mitigation strategy. Although it provides many benefits, it has significant problems that make its wide use infeasible. The multi-run mitigation strategy incurs heavy computational costs, which makes it infeasible in combination with complex models. Finally, instead of mitigating the variance, it only provides an approximation of the results distribution.
Mitigation strategies have the potential to have the highest impact on all the task for dealing with effects of randomness and on the use of learning with limited labelled data in practice. Having an effective mitigation strategy that can take the interactions into consideration will positively impact all open problems mentioned in this section. Therefore, it is important to focus on designing mitigation strategy that can address the problem as a whole. Existence of a single mitigation strategy that can efficiently mitigate effects of randomness from all randomness factors is probably infeasible. Therefore, the focus should be on effectively combining different randomness factor specific mitigation strategies. Besides taking interactions into consideration and achieving significant reduce in randomness, a specific focus should also be given to reducing the additional computation cost that is incurred. This will be especially important for the multi-run strategy, as we see it being used for randomness factors without specific strategies designed for them.
8.7. Limited Consideration For Randomness in Comparisons and Benchmarks
Even though many modifications are proposed that can deal with effects of randomness, they are applied only in specific cases of comparisons between approaches, but not as a part of tasks for addressing the effects of randomness. The results from experiments investigating the randomness factors and especially determining the effectiveness of mitigation strategies are compared in a simple manner based only on aggregated value. In case of mitigation strategies, it is common to even use a single run. Similarly to comparing approaches, the experiments investigating randomness are still sensitive to the effects of randomness as not all randomness factors are considered. More sophisticated comparison strategies should be applied when comparing between different randomness factors. Even though we can take inspiration from comparisons between approaches, not all of them are applicable.
In addition, the benchmarks are only starting to consider randomness in their design. Even when randomness is considered in the design, it is usually in a simple manner, such as looking at simple characteristic, providing multiple splits to train on, but the number of such splits provided is rather low (Zheng et al., 2022; Mukherjee et al., 2021). Even this little consideration is a step in good direction, but we need benchmarks that explicitly take variance into consideration and use it to penalise approaches that are brittle in presence of randomness, such as in (Efrat et al., 2022). Finally, some randomness factors are not even considered in benchmarks, such as model initialisation or the choice of data. Therefore, we still have a long way to go to have benchmarks that take into consideration all the important randomness factors.
9. Conclusion
In this paper, we have filled the gap of a missing, systematic literature survey focused on the existing research works dealing with effects of randomness that negatively impact the stability in real-world settings when labelled data are lacking. We summarised the current approaches for investigating the effects of randomness, determining the origin of randomness and designing mitigation strategies for different randomness factors. In addition, we analysed a specific group of papers that aim to take the randomness into consideration when comparing and benchmarking different machine-learning models. Based on the findings from the comprehensive analysis and synthesis of existing works, we identified a list of challenges and open problems when dealing with effects of randomness and provided some insights into how to handle them. We hope that this comprehensive and systematic literature survey will help researchers more effectively understand the tasks performed when dealing with the effects of randomness and grasp its core challenges so that the field can be advanced. In addition, we expect that it will encourage researchers to focus on the underrepresented randomness factors, their interactions and differences between their effects across different machine learning approaches so that the effects of randomness and the behaviour of models when they are present can be understood in more detail.
Acknowledgements.
This research was partially supported by TAILOR, a project funded by EU Horizon 2020 research and innovation programme under GA No. 952215; DisAI, a project funded by European Union under the Horizon Europe, GA No. 101079164; and by vera.ai project funded by the European Union under the Horizon Europe, GA No. 101070093.
Appendix A Detailed Paper Categorisation Using the Defined Taxonomy
As a part of the main content of the survey, only a part of the full paper categorisation of the core papers according to the defined taxonomy is included, mainly due to its size. The full paper categorisation, along with additional relevant information, such as a basic categorisation of the recognise papers is a part of digital appendix of this survey, which we plan to continuously update as new papers are released [^1] Available at https://docs.google.com/spreadsheets/d/1PW4PSq_mnfxkzVqxFw71XPqUumQ5GVJsc-Yfxa_yUdc.
Following information is part of the digital appendix:
  1. (1)
    Full categorisation of core papers – The categorisation is extended with additional taxonomy dimension as well as other metadata. The machine learning approaches are further divided into more detail, e.g., in meta-learning, we identify which papers deal with optimisation-based meta-learning and which consider metric-based meta-learning. In addition, we also divide the tasks performed to the full list presented in the survey to provide better overview of which tasks the papers focus on when addressing the effects of randomness. We also add the dataset secondary dimension that indicates what datasets the individual papers work with. Besides the changes to the taxonomy dimensions, we also include additional metadata, such as Document Object Identifier (DOI) number (where applicable), year of publication, or link to the paper.
  2. (2)
    Basic categorisation of recognise papers – We include the full list of the recognise papers in this digital appendix, along with their metadata (authors, DOI, year of publication). In addition, we identify some of the randomness factors the recognise papers deal with. Finally, we specify a high-level idea for each paper, for example, whether the paper recognises the problem of instability, tries to improve the situation.
  3. (3)
    Evolution in number of papers through years – Using the year of publication for each identified paper, we show the evolution in the number of papers that focus on addressing the effects of randomness.
  4. (4)
    Relation mapping between randomness factors and machine learning approaches for different tasks performed – As a part of the survey, such mapping was presented in aggregated form across all the tasks for addressing the effects of randomness. Besides the aggregated mapping, we also include the same table, but for each task separately. This mapping serves as an additional overview of what the individual papers focus on (randomness factors in specific machine learning approaches and tasks performed).
Appendix B Description of Machine Learning Approaches
In this part, we provide basic idea and high-level description of the different machine learning approaches that specify our scope in the survey, along with some representatives for each. This should serve as a basic overview for understating the concepts of the approaches required for better understanding of the surveyed papers, not as a comprehensive overview of the selected approaches. We provide description for meta-learning, language model fine-tuning and in-context learning approaches.
B.1. Meta-Learning
The idea behind meta-learning is to learn a general learning algorithm that can generalise across different tasks, with only few samples available for each task. In essence, meta-learning learns a function (also called meta-learner) that takes a set of training data as input and outputs the learned models parameters (also called base-learner) that can be used for the specific task. The few-shot learning setup is often used, where the samples are distributed across multiple tasks. During training, a subset of these tasks is sampled and split into training and testing set. Using the training tasks, the parameters of the base-learner are optimised. The parameters of meta-learner are then updated based on how good the base-learner performs on tasks in testing set. This training is performed over multiple iterations, while at the start of each iteration, the base-learner is initialised or updated based on the information or knowledge contained in the meta-learner. At test time, the base-learner is primed for the specific task using a few adaptation (or support) samples.
Different approaches of meta-learning differ based on the objective they optimise for and how the base-learner and meta-learner are defined. In optimisation-based meta-learning, the objective is to find an optimal set of parameters in the meta-learner that can be quickly adapted to any task when fine-tuned as a part of base-learner on a given task. In optimisation learning, the base-learner and meta-learner often represent the same model, and the base-learner is initialised by simply copying all the parameters of meta-learner. The optimisation of the meta-learner is performed similar to the supervised learning (where the error is back-propagated through the whole network), but using a second order derivative. Majority of optimisation-based approaches are based on Model-Agnostic Meta-Learning (MAML) (Finn et al., 2017), as they are designed to overcome some of the problems present in MAML, for example reducing its computational complexity by approximating the outer loop (e.g., Reptile (Nichol et al., 2018) or First Order MAML (Finn et al., 2017)) or optimising only part of the network while having the rest pre-trained (e.g. ANIL (Raghu et al., 2019)), or better dealing with task dissimilarity and distribution (e.g. MMAML (Vuorio et al., 2019), LEOPARD (Bansal et al., 2020)).
In metric-based meta-learning, the objective is to learn good representation for comparing adaptation samples with samples without known labels. The final prediction of the base-learner is determined by the class of the most similar adaptation (or support) sample. Most popular approaches for metric-based meta-learning are Prototypical Networks (Snell et al., 2017), or Matching Networks (Vinyals et al., 2016), which only differ in how they function (e.g., comparing based on aggregated information from multiple samples or comparing between each sample separately).
For a more comprehensive overview of taxonomy of different meta-learning approaches and more detailed description, please refer to (Hospedales et al., 2021; Lee et al., 2022b; Tian et al., 2022).
B.2. Language Model Fine-Tuning
The main idea of language model fine-tuning is to transfer knowledge from a pre-trained model to a specific task by updating its learned parameters using a few samples. The language model is first trained on a large corpora of unrelated texts, or an already pre-trained language model is used. The last layer of such language model is then replaced, or a new one is added for the specific task. Afterwards, the language model is fine-tuned on the few samples available by simply running a training process for few episodes on the samples with much lower learning rate.
As the process is pretty straightforward, the approaches differ only in what language model is used (which also defines what corpora it was pretrained on) and what part of the language model is fine-tuned. The common approaches include training all the layers, or only last few layers of the language model (such as only the added classification layer), or even approaches that fine-tune a specific, often disconnected layers (Lee et al., 2022a). The most popular language models that are used for fine-tuning are BERT (Kenton and Toutanova, 2019) and its modification like RoBERTa (Liu et al., 2019), ALBERT (Lan et al., 2019), DistilBERT (Sanh et al., 2019), XLNet (Yang et al., 2019) and many others.
For more comprehensive overview of different language model fine-tuning approaches, please refer to (Rogers et al., 2020; Xia et al., 2020),
B.3. In-Context Learning
In-context learning is an emerging paradigm, where a large pre-trained language model is used for different tasks without first updating its parameters. Instead, all tasks are reformulated as a sequence generation problem and the language model is “prompted” to output a sequence of words. The resulting sequence of words is then mapped to a possible list of labels (e.g., word ”good” is mapped to positive and ”terrible” to negative label in sentiment classification). To prime the model for a specific task, the presented input is constructed as a concatenation of a few labelled samples (serving as context) and a single test sample, for which the label is predicted.
As the final prompt is a concatenation of multiple labelled samples and single unlabelled one, usually only language models that allow large input sizes can be used. Another important choice in in-context learning is how the input prompt is designed, either manually, semi-automatically or automatically. The approaches differ on what language models are used, how the prompt format is designed and whether the in-context learning is combined with some kind of fine-tuning. The most popular model used in in-context learning is GPT-3, with the approaches starting to design automatically generated prompts, although also smaller models are used when combined with fine-tuning (Brown et al., 2020; Gao et al., 2021; Perez et al., 2021; Logan IV et al., 2022).
For more comprehensive overview of different in-context learning approaches, please refer to (Liu et al., 2023; Dong et al., 2022; Song et al., 2023).
Appendix C Implementation of the Survey Methodology
C.1. Search query definition
The search query is composed of two parts: 1) terms that that represent the all the versions of stability that we consider in this survey; and 2) scope describing our focus on approaches that can be used for learning with limited labelled data.
The set of terms that can be used to describe the effects of randomness on the performance was progressively expanded based on the terms in papers that in any way address this randomness. The identification of the papers were done randomly by first sampling papers that reference stability, assessing whether their understanding of stability adheres to our definition, and using the cited and citing papers to identify next potential candidates. Using this methodology we identified first set of words that included following words: stability, instability, sensitivity, variance, randomness, robustness and reproducibility. However, we decided to remove the terms robustness and reproducibility from this list, as they have much wider meaning and would lead to many irrelevant papers.
The scope was defined based on groups of approaches we want to focus on. Based on reading the papers, we have noticed only two terms being widely used to define all the approaches we were interested in: meta-learning and few-shot learning. We decided to use the term few-shot learning instead of learning with limited labelled data, because of its popularity and its misuse leading to both term meaning the same thing in the literature. We noticed many different uses of these terms across papers, with many removing the hyphenation or foregoing the ”learning” part of few-shot learning. To account for all the different uses of the terms, we finally decided to use the most simplified version of them, specifically meta learning, metalearning and few shot.
The final query we used for the keyword search was: ‘(“few shot” OR “meta learning” OR “metalearning”) AND (“stability” OR “instability” OR “sensitivity” OR “variance” OR “randomness”)‘. We also searched only for papers that were published in the year 2017 or later, as the focus on effects of randomness and stability before this year was non-existent.
C.2. Relevant Digital Libraries Used to Discover Papers
We used the digital libraries that provide flexible search options, with focus on machine learning domain as one of its domain and allow for reproducible search results to discover the relevant papers. As we were not interested in other disciplines, we limited the search results to only those relevant for us. This includes following databases: ACM Digital Library, IEEE Xplore and Scopus database.
In addition, we also searched through the proceedings of the top rated machine learning conferences, such as ICML, ICLR, or NeurIPS. We have noticed that many relevant papers are submitted to these conferences, while at the same time only a fraction of them is indexed by the previously stated databases. We specifically searched through ACL Anthology (containing the aggregated proceedings of the ACL conferences), the NeurIPS conference and the Proceedings of Machine Learning Research (PMLR), which includes proceedings from conferences such as ICML or ICLR. Only the ACL anthology allows for advanced search through papers, nevertheless, the search is limited to only top 100 results. Therefore, to search through these libraries we opted to use Google Scholar. Each conference was searched separately using a modified query to limit results to only specific site (using the keyword in search, i.e. ‘site:aclanthology.org‘ for ACL, ‘site:neurips.cc‘ for NeurIPS and ‘site:mlr.press‘ for PMLR). We take only the top 200 relevant papers from each conference. To determine the cut-off point, we manually determined the relevance of the papers at later pages of results and noticed the relevance drops of significantly after $\sim$100 papers. Therefore we decided to search through 100 more papers just to be sure we are not missing any relevant paper.
Lastly, we also use included cited and citing papers from the most relevant ones as additional sources. We detected a significant clustering behaviour in papers that deal with the effects of randomness, with papers related to specific groups of approaches extensively citing all other papers in this group. This significantly improves the coverage of results as it includes papers not indexed in the databases, but also those that may use slightly different terminology.
C.3. Identification of relevant papers using search and filtering
Since the research paper collection from digital libraries would naturally result into a large set of papers with various relevance, we selected the subset of the most relevant papers that we call core papers. To identify the set of these papers, we perform following steps across multiple iterations:
  1. (1)
    Create set of potentially relevant papers. The defined search query is used across all specifies digital libraries to create the set of papers that are potentially relevant for the work. This generates a large set of papers with many false positives.
  2. (2)
    Filter papers based on relevance. The large set of potential papers is reduced to a set of papers dealing to some extent with the effects of randomness. To filter out the irrelevant papers the context in which the individual key words from the search query are used in the paper. This is done by finding all the appearances of the key words using full-text search and manually checking them. After finding first relevant context for both the term and scope key words, the paper is considered to be relevant. If no such context exists for one of the groups, or the context indicates different definition of the problem (such as algorithmic stability), the paper is considered to be irrelevant and removed from the set. For example, we do not consider the context to be relevant if it appears in related work. This step significantly reduces the set, leaving only 5̃of papers.
  3. (3)
    Filter papers based on merit. Although the papers in this step are mostly relevant, the set still contains papers we do not want to consider based on their form. This includes papers that are only extended abstracts, proposals for talks or projects, or papers appearing on papers not relevant to stability. In addition, papers from conferences with lower Core rank than B or from journals not belonging to Q2 or Q1 are also removed.
  4. (4)
    Filter papers based on focus. In this step we filter out the papers that only recognise that there is a problem with stability in regards to randomness in some approaches, but provide no further investigation or focus on the randomness. We call this papers as recognise papers. This mostly includes papers that note that the training process with limited data is unstable and so they perform multiple runs, or papers that note that it was found in previous papers that there is a problem. This leaves us with set of papers we consider to be core.
  5. (5)
    Extend set of papers with citing and cited papers. As there is a strong clustering present, papers dealing with similar approach cite each other extensively. We make use of this clustering to discover other relevant papers that may not be indexed by the databases or that use different terminology and so the search query does not catch them. However, we extract such papers only based on the context they are referenced in (when cited in the paper), or based on a the relevant filtering step applied only to the title and abstract of the paper (which cites the current paper). We identify the citing papers through the Google Scholar cited by section present next to each paper.
We repeat the steps $2-5$ until no new papers are identified in step $5$.
C.4. Analysis and categorisation of the papers
All the papers identified this way were then manually categorised using our taxonomy as described in the main content of the survey (Section 3). The content of each paper was analysed and assigned value for each defined property. Afterwards, the main categorisation property (task performed to address the randomness) was selected based on how best it splits the papers into independent groups.
For each task for addressing the effects of randomness, additional properties and characteristics were identified and used to cluster the papers further. These additional properties were then used to inform the proposal of the main findings from each tasks. Therefore, each finding can be mapped to one or multiple such properties at the same time.
The challenges and open problems were identified by observing common patterns in the findings, such as contradictions in results. In addition, the open problems were further extended by identifying parts of the research process that are currently missing even though their inclusion is a straightforward and simple modification of the current approach, such as using statistical evaluation of results instead of reporting a single performance metric.
dWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">